### v3

# Analyzing the Effect of Linguistic Similarity on Cross-Lingual Transfer: Tasks and Experimental Setups Matter

Abstract
Cross-lingual transfer is a popular approach to increase the amount of training data for NLP tasks in a low-resource context. However, the best strategy to decide which cross-lingual data to include is unclear. Prior research often focuses on a small set of languages from a few language families and/or a single task. It is still an open question how these findings extend to a wider variety of languages and tasks. In this work, we analyze cross-lingual transfer for 263 languages from a wide variety of language families. Moreover, we include three popular NLP tasks: POS tagging, dependency parsing, and topic classification. Our findings indicate that the effect of linguistic similarity on transfer performance depends on a range of factors: the NLP task, the (mono- or multilingual) input representations, and the definition of linguistic similarity.
Analyzing the Effect of Linguistic Similarity on Cross-Lingual Transfer:
Tasks and Experimental Setups Matter
Verena Blaschke1,2,*   Masha Fedzechkina3   Maartje ter Hoeve3 1Center for Information and Language Processing (CIS), LMU Munich 2Munich Center for Machine Learning 3Apple blaschke@cis.lmu.de, {mfedzechkina, m_terhoeve}@apple.com
*Work done while interning at Apple.
1 Introduction
For many of the world’s languages, the available data to train natural language processing (NLP) models is scarce. If data is available, it is often only enough for an evaluation set, raising the question of how to select the training set. Two approaches are intuitive: (i) based on linguistic similarity measures: find the training data in a language that is linguistically as close as possible to the target language, where “linguistically close” is defined by one or multiple linguistic similarity measures, or (ii) based on dataset dependent measures: find a dataset in another language that is similar to the target dataset, e.g., in terms of high word or $n$-gram overlap. Naturally, these two types of approaches are not mutually exclusive.
Figure 1: Languages included in our experiments. Green indicates languages included in all tasks, blue languages only used for POS tagging and parsing, and purple languages only used for topic classification. Base map via naturalearthdata.com (CC0).
However, it is unclear which of these measures are most important to select the source language for cross-lingual transfer, despite previous work focusing on this question (see §2): earlier studies often contradict each other, and therefore leave a number of important avenues for improvement. For example, prior work often lacks a large representation of languages and language families, as well as NLP tasks, and sometimes relies on synthetically constructed datasets. Furthermore, popular out-of-the-box measures for linguistic similarity are sometimes intransparent or faulty (Toossi et al., 2024; Khan et al., 2025; §3.3), and correlations between different similarity measures are often not taken into account. Summarizing, how findings from prior work generalize across a larger variety of languages and tasks, and how we should interpret different similarity scores, are open questions.
In this work, we contribute to these questions by analyzing transfer between 263 different languages from 33 language families (Figure 1) in three different tasks: (i) part-of-speech (POS) tagging, (ii) dependency parsing, and (iii) topic classification. In choosing our tasks, we are motivated by the following considerations: (i) availability of datasets: we select tasks for which we can find data for a multitude of languages from different language families, and (ii) types of tasks: we include two word-level grammatical tasks (i.e., POS tagging and dependency parsing) and one sentence-level topic identification task. To allow for a clean analysis, we opt for a zero-shot approach, in line with prior work de Vries et al. (2022) – we train our models on the task in one particular source language, and evaluate on the target language without additional training or fine-tuning in the target language.
Our key findings can be summarized as:
  1. 1.
    For different tasks and input representations, different similarity measures are most predictive for cross-lingual transfer performance. For instance, syntactic similarity is most predictive for POS tagging and parsing, whereas trigram overlap is the most important predictor for $n$-gram-based topic classification (§5.1);
  2. 2.
    Practical implication within studied tasks: Choosing a source language based on a pertinent similarity measure leads to adequate transfer results (§5.2.1);
  3. 3.
    Practical implication across studied tasks: When no information about pertinent similarity measures is available for a given task, choosing a source language based on findings for a conceptually similar experiment is a relatively safe choice (§5.2.2).
2 Related Work
Table 1 provides an overview of related work, highlighting an apparent trade-off between including many languages or including multiple tasks. An important difference that distinguishes our work from prior work is the focus on both. We argue that including a large number of both source and target languages is important for ensuring a relatively balanced distribution of languages. [^1] I.e., such that there is not mainly one cluster of (European) languages that are similar to each other, and most other languages are both dissimilar to that cluster and to each other. In comparison to the studies that come closest to our work in terms of number of included source and target languages de Vries et al. (2022); Samardžić et al. (2022), we include a larger variety of tasks, allowing us to draw conclusions across task boundaries.
Philippy et al. (2023) survey contributing factors for cross-lingual transfer, including most of the works in Table 1. An important take-away from their work is that prior research presents contradicting findings. E.g., some studies find lexical overlap to correlate more strongly with token rather than sentence classification tasks Srinivasan et al. (2021), whereas others find the opposite Ahuja et al. (2022). Similar contradictory examples are presented for different linguistic similarity metrics, pre-training dataset size, and model architecture.
Based on this insight, Philippy et al. make recommendations for follow-up work: (i) focus on real natural languages (instead of synthetic ones), (ii) examine the interaction between different contributing factors, (iii) focus on many languages, (iv) focus on linguistic features when selecting training languages, and (v) focus on generative tasks, given the success of generative models. We focus on the first four recommendations in this work. We include three classification tasks (POS tagging, dependency parsing, and topic classification), motivated by the availability of train and test data in a large number of languages for these tasks.
[…]
Table 1: Related work focusing on zero-shot transfer between many languages or on many tasks. Tasks: D=dependency parsing, E=entity linking, F=natural language inference, I=intent classification, M=machine translation, N=named entity recognition, P=part-of-speech tagging, Q=question answering, S=sentence retrieval, T=topic classification.
3 Methodology
We run transfer experiments on two word-level grammatical tasks with word-level annotations (§3.1) and a sentence-level topic classification task (§3.2). Of the 263 languages in our experiments, 55 training and 84 test languages are shared between all tasks. Appendix §A contains details on all languages and datasets in this study. Section §3.3 introduces our similarity measures.
3.1 Grammatical tasks
Data
We use POS tags and syntactic dependencies from Universal Dependencies (UD; de Marneffe et al., 2021). The advantage of UD is the large number of languages (153) and language families (28) in which manually annotated datasets are available that both showcase a language’s specific syntactic structures and adhere to a shared set of annotation guidelines. UD also has its drawbacks in that the different treebanks are not necessarily from the same domains, most treebanks were annotated independently by different groups of researchers, and the size of the train and test splits is not identical across treebanks. To mitigate the latter, we evaluate the effect of training dataset size in §5.1. To account for differences in the test sets (e.g., differing complexities or sentence lengths), we focus on comparing the performance of different parsers/taggers on each test set (rather than comparing how well each parser/tagger does across test sets; §5.1).
We use the test splits of UD release (2.14; Zeman et al., 2024): 268 treebanks in 153 target languages, from 28 language families. We use pretrained models (see below), trained on 124 treebanks in 70 source languages from 12 language families. We exclude treebanks that are without text, particularly small, focus on code-switching, or have inconsistent train/test splits (§A.1).
Models
We use the UDPipe 2 models (Straka, 2018, 2023) that were trained on UD 2.12 data. Each model is trained on a single treebank. The models have so far only been systematically evaluated on their within-treebank performance, but not cross-lingually. UDPipe 2 combines monolingual character and word embeddings with multilingual embeddings derived from mBERT Devlin et al. (2019). The models are trained jointly for POS tagging, dependency parsing, lemmatization and morphological feature prediction. We only evaluate on the first two, as many treebanks do not have gold standard labels for the others. UDPipe 2 post-processes the predicted dependencies to ensure that each sentence includes a root dependency that all other nodes (in)directly depend on.
We evaluate POS tagging using accuracy, and dependency parsing using the labeled (LAS) and unlabeled attachment scores (UAS). For LAS, we follow UD’s evaluation scripts and ignore dependency label subtypes.
3.2 Topic classification
Data
We use the SIB-200 dataset (Adelani et al., 2024), a subset of FLORES-200 (NLLB Team et al., 2022) with parallel sentences in 194 languages [^2] This excludes three Arabic varieties whose sentences are nearly identical to the Standard Arabic sentences (§A.2). (from 22 families) annotated with seven topic labels. Eight languages are represented twice, but with different writing systems. SIB-200 contains 701 training and 204 test sentences.
Models
We use multi-layer perceptrons (MLPs) for topic identification, similar to the baseline models by Adelani et al. (2024). This lightweight architecture allows training and evaluating many models without prohibitive time or energy investments, and results in evaluation scores that are close to the performance of base-sized transformers like XLM-R Conneau et al. (2020)D, Table 2). We use the scikit-learn implementation (Pedregosa et al., 2011) and conduct hyperparameter tuning on a subset of the languages (details in Appendix §C). We compare different ways of representing the input data:
  1. 1.
    Character $n$-gram counts (topics-base). We use character-level $n$-gram counts ($1$- to $4$-grams) to represent the input. This ensures that we know exactly what training data were used, and it puts all languages on equal footing.
  2. 2.
    Transliterated input (topics-translit). To remove differences between writing systems, we use uroman Hermjakob et al. (2018) to transliterate the dataset into Latin characters and remove diacritics, and otherwise repeat the previous set-up. We exclude three datasets that were not supported by uroman (§A.2).
  3. 3.
    Multilingual representations (topics-mbert). To allow for a more direct comparison with the grammatical experiments, which partially rely on multilingual mBERT representations, we follow UDPipe 2 by deriving embeddings from the mean-pooled last four layers of the frozen base-sized, uncased mBERT model Devlin et al. (2019). We use the hidden representations of the [CLS] token as input to the MLP.
3.3 Similarity measures
We include a range of dataset-dependent and -independent similarity measures, which are similar to measures used in related work (§2).
Structural similarities
Grambank Skirgård et al. (2023) encodes grammatical information for several thousand languages. Its 195 grammatical features are chosen to allow almost no logical dependencies between the values of different features, i.e. the value of one feature does not logically entail the value of another.
We additionally use the lang2vec tool (Littell et al., 2017, 2019), which has also been used in many other studies on multilingual NLP (cf. Toossi et al., 2024). Lang2vec aggregates information on syntax, phonology, and phoneme inventories from multiple databases in the form of binary features. Since not all sources contain full information on all languages, we use language vectors that include information from multiple of the above sources (averaged values where sources disagree). Some grammatical features are covered by both Grambank and lang2vec, although sometimes with different value assignments Baylor et al. (2023).
We use Gower’s (1971) coefficient to calculate similarities between language pairs by comparing their feature values. If information is missing for a feature in one or both languages, we ignore that feature. If two values are identical, their similarity is 1, otherwise it is 0. The overall distance between the two languages is the mean of the (attested) feature distances. We calculate similarity scores for Grambank’s entries (gb) and for lang2vec’s syntactic (syn), phonological (pho), and phonetic (inv) entries. We include only similarities for language pairs where at least half of all features could be included in the calculations. However, in practice, more features are compared in most cases. On average, 83% of the features are included in each similarity calculation for gb, 63% for syn, 100% for inv, and 84% for pho.
Lexical similarity
As a proxy for how similar different languages’ vocabularies are, we compare multilingual word lists from the Automated Similarity Judgment Program (ASJP; Wichmann et al., 2016). Jäger (2018a, b) calculated language dissimilarity scores based on ASJP, taking into account cross-linguistic phonological patterns. Our lexical similarity score (lex) is 1 minus Jäger’s dissimilarity score. Since some languages have multiple ASJP entries, we use language-wise mean scores.
Phylogenetic relatedness
We determine whether two languages are related (and how closely) with the help of Glottolog’s Hammarström et al. (2024) phylogenetic trees. For each language, we retrieve the path from its family root node to the language node. The relatedness of two languages (gen) is the ratio of shared nodes along these paths. [^3] Because different trees/branches have different depths (owing to different family sizes and/or different levels of documentation), comparisons of scores can only serve as a proxy for the degree of phylogenetic relatedness. This is also noted in the metadata for lang2vec, which calculates genetic distances in a similar way: http://www.cs.cmu.edu/~dmortens/projects/7_project/. For constructed languages, we ignore the top-most level (e.g., “artificial languages”) and instead retrieve phylogenetic information starting at the second level.
Geographic proximity
We use lang2vec’s location information (each language is represented as a vector of distances to a number of points on the Earth’s surface) and calculate the Euclidean distances between language vectors. [^4] We do not use lang2vec’s precomputed distances, as the package erroneously returns the maximum distance for some languages, regardless of the comparison. We define geographic proximity (geo) as 1 minus the distance.
Character and word overlap
We measure the overlap between training and test datasets on the character level (chr) with the Jaccard similarity of the character sets. We repeat this on the word level (wor) for the UD datasets (where the data come with gold-standard word tokenization), and on the character trigram (tri) and mBERT subword token (swt) level for SIB-200.
Amount of training data
We count the number of sentences in each training dataset (size). For the topic classification task, this is trivial as each language has the same number of training samples.
Correlations between measures
Correlations between similarity measures can influence how importance is assigned to different measures in regression analyses. Table 8 in §B shows how the different similarity measures are correlated with each other. Lexical similarity and phylogenetic relatedness are highly correlated ($r=0.87$). Additionally, the two grammar-based similarity measures (gb and syn) are moderately strongly correlated ($r=0.61$), as are gb and lexical/phylogenetic similarity ($r=0.57$, $0.56$), and word overlap in UD and lexical similarity ($r=0.59$).
4 Supporting Results
Here we present the results on which the main analyses in §5 build. In §4.1, we present the POS tagging, parsing and topic classification scores within and across languages. In §4.2, we compare how (dis)similar transfer patterns are across tasks and experiments, which helps contextualize the importance of different similarity measures for different experiments and tasks in §5.1.
4.1 Within-language vs. cross-lingual performance per task
As expected, the within-language performances are much higher than the cross-lingual performances. Table 2 shows the mean scores within and across languages and datasets. In the following analyses, we focus on the language-level results. For UD, which has nearly twice as many datasets as languages, we find that trends are very stable across datasets of the same language (§D.3).
All of our models achieve reasonable within-language results, indicating that cross-lingual transfer (to an appropriate target language) could be feasible for all models. For the two straightforward classification tasks (POS tagging and topic classification), we construct random baselines (§D.1) that are outperformed by all models in within-language evaluations. For our topic classification models, the within-dataset performance is comparable to the baselines in the dataset paper (Adelani et al., 2024), but we do not match the performance of their best model. Our POS tagging results are similar but not identical to the ones that de Vries et al. (2022) obtained on UD v2.8 with fine-tuned XLM-R models for 65 training and 105 test languages. If we consider our own POS tagging results but only select the subset of languages that was used in de de Vries et al.’s experiments, the POS tagging accuracies are highly correlated with those that de Vries et al. obtained (Pearson’s $r$ and Spearman’s $\rho=0.73$, $p<0.0001$). Thus while the model choice (UDPipe vs. XLM-R) matters for the resulting patterns, many transfer trends are similar.
Dataset level Language level Within Across Within  Across Grammatical tasks POS accuracy 96.4 3.1 43.6 20.8 93.4 6.8 43.0 20.2 POS acc. (de Vries et al.) 94.1* 4.5 57.4* 22.4 94.1 4.5 57.4 22.4 UAS 88.7 6.4 37.3 19.3 84.8 10.1 36.8 18.8 LAS 84.6 8.6 21.2 17.9 78.4 13.7 20.6 17.0 Topic classification (accuracy) topics-base 70.3 4.0 20.7 8.8 66.7 14.1 20.1 8.8 topics-translit 69.4 4.1 22.5 8.2 66.7 10.8 22.5 8.2 topics-mbert 60.1 18.7 42.6 20.4 58.2 19.9 42.6 20.4 MLP (Adelani et al.) 62.3 XLM-RB (Adelani et al.) 70.9 XLM-RL (Adelani et al.) 76.1
Table 2: Scores of models evaluation within vs. across datasets and languages, in this work and related work. Our analyses focus on the language-level scores. Scores are mean scores (in %), with standard deviations in subscripts. XLM-RB and L = XLM-R base/large. *When multiple datasets were available for one language, de Vries et al. combined them into one.
For topic classification, we compare the different input representations. The within-language performance is higher for the monolingual $n$-gram-based models than the models with multilingual input representations from mBERT. We hypothesize that this due to high [UNK] token rates for some languages that were not in mBERT’s pre-training data. [^5] Bagheri Nezhad and Agrawal (2024) similarly find that the performance of multilingual models that are monolingually trained and evaluated on SIB-200 relates to how well a language is represented in a model’s pre-training data.
Although topics-base and topics-translit achieve the same within-language accuracy, topics-translit performs slightly better cross-lingually, having the advantage of higher $n$-gram overlap between datasets. [^6] We further analyze the effect of writing systems in §D.4. In our correlation analysis (§5.1), (lack of) character overlap serves as a proxy for writing system differences. In cross-lingual evaluations, topics-mbert benefits from its multilingual input embeddings: its accuracy is about twice as high (42.6%) as for the monolingual models. Generally, transfer works well if both training and test languages are in mBERT’s pre-training data (or closely related to a language that is): The average accuracy is 68.3% if both languages were included in mBERT’s pre-training data, 46.5% if only the target language was included, 36.1% if only the source language was included, and 33.2% if neither was included. This is consistent with results by Adelani et al. (2024), who fine-tuned four XLM-Rlarge models on one high-resource language each and found their transfer results to be very similar to each other.
POS accuracy     Dependency UAS   Dependency LAS .
Refer to caption Refer to caption Refer to caption
topics-base        topics-translit topics-mbert     .
Refer to caption Refer to caption Refer to caption
Figure 2: Different experiments produce different transfer patterns. NLP transfer results for all combinations of training (columns) and test languages (rows). The darker a cell, the higher the score. The three heatmaps for the grammatical tasks are sorted in the same order, and the three heatmaps with the topic classification results are sorted in the same order. The darker diagonal shows the within-language scores. Large, labelled heatmaps are in Appendix §D.2.
4.2 Comparing transfer patterns across tasks
In Figure 2, we plot heatmaps of the transfer results for the different experiments. Darker colours correspond to higher scores. To save space, we omit the language labels (which can be found in §D.2). We compute correlations between the results, which we summarize below (details in §E.1).
The results of the grammatical tasks are highly correlated with each other, the correlations between other task results are weaker. [^7] Because of the very high correlation between the parsing measures LAS and UAS (Pearson’s $r=0.95$), we focus on LAS in the remainder of the paper as the more extensive of the two: LAS not only measures whether the dependency arcs are correct, but also whether they were labelled correctly. Even when using exactly the same underlying text data and models, evaluating on a related task results in transfer trends that are similar, but not entirely the same: The parsing (LAS) and POS tagging results by identical UDPipe models (except for the classification heads) are correlated with $r=0.86$.
The preprocessing of the input data (e.g., transliteration) and especially the choice of input representation (mono- vs. multilingual) affects the transfer trends: The results of the two $n$-gram-based topic classifiers are highly correlated with each other ($r=0.68$), but not with the results of the mBERT-based set-up ($r=0.28$ and $0.36$). Set-ups that involve multilingual representations are more highly correlated. For the languages that appear both in UD and SIB-200, the results of the grammatical tasks are most strongly correlated with those of the mBERT-based topic classification models ($r=0.64$ for POS tagging, $r=0.58$ for LAS).
5 Main Results and Analysis
Here, we investigate which factors correlate with overall transfer performance (§5.1). Then, we explore what this means for selecting a source language for cross-lingual transfer (§5.2).
5.1 How do the similarity measures correlate with the tasks and input representations?
We calculate the correlations between the similarity measures and the NLP task results. To account for differences between test sets, we calculate correlations (Pearson’s r) between similarity measures and transfer results on a test language level. We compare the performance of different NLP systems on the same test language, but we do not compare the performance of a single system across multiple test sets. This decision is motivated especially by the challenges of comparing parsing performance across treebanks: Sentence length influences parsing difficulty as it determines the available search space (cf. McDonald and Nivre, 2011, Choi et al., 2015), and morphological differences between languages hamper the comparability of parsers Nivre and Fang (2017).
We analyze overall correlation scores by averaging across test languages. We treat correlation coefficients with p-values of at least 0.05 as 0, and exclude items where we could not calculate similarity scores due to missing linguistic information. We use language-level averages so that those languages with multiple training and test datasets do not have artificially high correlations, and languages with multiple test datasets do not have greater influence on the overall averages. However, if we instead consider treebank-level correlations, we observe that test sets that belong to the same language show very similar correlation patterns and scores (§D.3).
The correlations between task results and similarity measures vary across our experiments. Figure 3 shows the correlations, which we summarize below. Unaggregated correlations and 95% confidence intervals of the mean scores can be found in §E.3. These unaggregated correlations show that, while there are some outliers, the overall trends we report below accurately reflect the language-level trends.
Figure 3: Mean correlation scores between task results and similarity measures. “Word*” = overlap between words (wor; UD tasks) trigrams (tri; topics-base/translit), and subword tokens (swt; topics-mbert). Dotted lines are added for intelligibility.
We observe some common trends across experiments: Training dataset size does not matter much: it is the same across all topic classification experiments, and the correlations between training set size and performance in the grammatical tasks are close to zero. The phonological and phonetic similarity measures (pho, inv) also generally have low correlation scores. [^8] The correlations with inv are comparatively high for the $n$-gram-based topic classification models. We assume that this is due to inv being moderately correlated with string similarity measures like trigram overlap (§B), which is in turn strongly correlated with the topic classification results.
The strongest predictor for parsing performance is syntactic similarity (syn) as determined by lang2vec ($r$avg_syn$=0.57$), followed by the similarity of Grambank features (gb; $r$avg_gb$=0.42$). These correlations are even stronger when we only consider the test languages that mBERT was pretrained on ($r$avg_syn$=0.69$ and $r$avg_gb$=0.60$, respectively; Appendix §E.3).
The POS tagging outputs show correlation patterns similar to the ones for parsing, albeit weaker. Although the correlation strengths are generally weaker for POS tagging, e.g., the strongest predictor is lang2vec’s syntactic similarity (syn) with $r$avg_syn$=0.37$, word overlap (wor) is as good a predictor for POS accuracy as for LAS ($r$avg_wor=0.33 and $r$avg_wor=0.34). Again, the correlation strengths are higher when only considering test languages that were in mBERT’s pre-training data ($r$avg_syn$=0.46$). [^9] The correlation strengths are also higher when we analyze de Vries et al.’s results ($r$avg_syn$=0.46$). We hypothesize that this is due to the smaller amount of test languages in their experiments, a higher proportion of which is in XLM-R’s pre-training data. The correlations we observe for the grammatical task’s transfer results partially align with prior research: Lauscher et al. (2020), Samardžić et al. (2022) and Pires et al. (2019) also find syntactic similarity to be important, and Lin et al. (2019) and Xia et al. (2020) also find word overlap to be a good predictor for POS tagging. However, contrary to our results, the latter two find syntactic similarity to be relatively unimportant for both tasks. We hypothesize that such inconsistencies are due to differences in the language sets studied and other experimental differences (e.g., Lin et al. (2019) add data from the test language to the training set when possible).
The results of the $n$-gram-based models are most highly correlated with measures of string similarity and lexical similarity. This is expected based on their input representations. For topics-base, the trigram overlap (tri) between training and test data shows the highest correlation ($r$avg_tri=0.65), followed by character overlap (chr; $r$avg_chr=0.49) and lexical similarity (lex; $r$avg_lex=0.49). Trigram overlap and lexical similarity (and the highly correlated genetic relatedness, gen) are also the strongest predictors for topics-translit ($r$avg_tri=0.55, $r$avg_lex=0.61, $r$avg_gen=0.55). However, for this model, character overlap only plays a very small role ($r$avg_chr=0.11), as all datasets use the same character inventory due to the transliteration.
For the model using mBERT representations (topics-mbert), none of the correlations are strong. The correlations peak at $r$avg_chr=0.27 for character overlap (chr). Model performance depends instead on the inclusion of the source and especially target languages in mBERT’s pre-training data (§4.1).
The transfer results cannot be predicted by any one factor alone. We fit a linear mixed effects model for each experiment, with the NLP score as the dependent variable and the source and target languages as random effects. The fixed effects are the similarity scores, whether the training and test data use the same writing systems, and whether the test language was included in mBERT’s pre-training data. We use R R Core Team (2024) and the lme4 package Bates et al. (2015). This analysis involves a reduced set of languages, as for many language pairs in our data at least one similarity metric is not defined and we do not perform any data imputation. The analysis shows similar trends to the correlations described above. Additionally, for each experiment, most of the variables in our analyses are significant predictors of the transfer score (details in §E.2). This even sometimes applies to measures that capture similarity at the same linguistic level: e.g., both grammatical similarity measures (syn, gb) are independently among the highest predictors for parsing performance. The finding that multiple measures contribute to predicting the transfer score confirms previous work Lin et al. (2019); de Vries et al. (2022).
5.2 Practical implications
So far we have investigated correlations on a global level. In this section we investigate how these findings can be used in practice: Do the overall correlation patterns also apply when picking a source language for a given target language according to simple heuristics derived from the global patterns?
We intentionally pick very simple heuristics, as we believe them to be more realistic to how practitioners choose source language candidates in practice. Additionally, we choose them to be easily generalizable and not be constrained to the languages in our experiments. [^10] For instance, while LangRank Lin et al. (2019) is an easy-to-use tool for transfer recommendations, it is limited to the tasks and languages it was trained on and thus recommends one of 30 possible source languages for dependency parsing or one of 60 for POS tagging.
size pho inv geo syn gb gen lex chr word* Top-1 source candidate (= most similar language) POS deV 88 1011 1011 78 57 811 68 POS 2912 1513 1412 1515 1012 1212 0911 1010 1514 1213 LAS 2114 1312 1313 1316 710 1009 810 809 1616 1113 UAS 2712 1613 1512 1615 809 1311 910 1011 1714 1415 top.-b. 1712 1714 1314 1512 1413 912 911 1311 405 top.-tr. 1310 1311 $1$ 1109 1009 709 708 2013 304 top.-m. 1213 1113 1011 910 809 809 809 1210 913 Top-3 source candidates POS deV 34 45 45 25 24 58 24 POS 2513 708 708 507 304 406 507 406 708 507 LAS 1814 809 709 508 304 406 406 305 811 609 UAS 2413 808 708 608 305 506 507 507 910 709 top.-b. 1010 911 509 607 609 507 306 809 204 top.-tr. 808 709 408 506 506 305 305 1411 103 top.-m. 607 505 405 404 405 404 405 504 611
Table 3: Mean performance loss if picking source languages based solely on a given metric (or based on training dataset size). Performance loss in percentage points. Subscripts = standard deviations; “word*” = wor, tri, swt; deV = de Vries et al. (2022).
5.2.1 Should I pick the most similar language according to one similarity measure?
For each target language and each similarity measure, we calculate the difference between the best score obtained by any source language and the score obtained by the most similar language per the measure. We only consider the source language candidates for which we can compute a similarity score for the target. For size, we select the language with the most training data. The patterns for the most strongly correlated similarity measures are similar – choosing a source language based on the pertinent similarity measure incurs relatively small losses. E.g., parsing performance is most strongly correlated with syntactic similarity, and picking the syntactically most similar language as the source also results in the lowest performance loss for parsing (Table 3, top). Our findings suggest choosing a syntactically similar language for POS tagging and dependency parsing, and a dataset with high trigram overlap for the $n$-gram-based topic classification experiments.
Figure 4: Left: Relationship between phylogenetic and syntactic similarity – unrelated or distantly related languages can be syntactically similar or dissimilar, but all closely related languages are syntactically similar. Right: Relationship between character overlap (between training and test sets) and the classification scores of the topics-base model – transfer between languages with low character overlap works poorly, but high overlap does not guarantee good transfer.
However, some of the less strong predictors in the global correlations are nearly as good for picking a source language, e.g., genetic and lexical similarity for parsing. We hypothesize that this is due some linguistic measures being more correlated when similarity is high. [^11] E.g., very closely related languages tend to be syntactically similar, even if the overall correlation between genetic and syntactic similarity is only moderately high due the existence of languages that are unrelated but share syntactic similarities or that are (more distantly) related and syntactically dissimilar (Figure 4, left). Conversely, character overlap is a worse measure for selecting a source language than the overall correlations would suggest. This is likely due to transfer between languages with low character overlap generally working poorly, while high overlap does not guarantee good transfer (e.g., Figure 4, right).
We additionally simulate a setup where a researcher has data in multiple source language candidates at their disposal and can afford training and comparing a small selection of models. We select the top three most similar languages according to each measure and take the highest transfer score produced by any of them (Table 3, bottom). The same trends still hold as when picking only one candidate. However, the overall results are much better, and the gaps between the performance losses between the different measures become smaller. Comparing the results of multiple top source language candidates often yields better results than only considering the most similar one.
POS deV POS LAS UAS to.-b. to.-t. to.-m. Top-1 source candidate (= best src in another experiment) POS deV 304 406 506 610 712 606 POS 609 205 305 911 1012 1614 LAS 709 102 102 1114 1013 1814 UAS 810 305 102 1215 1012 1714 top.-b. 1211 1011 1112 1313 307 1915 top.-tr. 1009 808 808 909 306 1813 top.-m. 506 403 505 505 707 607 Top-3 source candidates POS deV 103 204 304 305 203 404 POS 306 102 103 406 406 707 LAS 305 001 508 508 910 UAS 406 103 001 508 507 809 top.-b. 608 507 508 508 103 1211 top.-tr. 506 506 405 506 205 1210 top.-m. 202 202 202 202 303 304
Table 4: Mean performance loss if picking source languages based solely on performance on the task in the column. Performance loss in percentage points. Subscripts = standard deviations; “word*” = wor, tri, swt; deV = de Vries et al. (2022).
5.2.2 Should I choose a source language based on another experiment?
We repeat these analyses but select source languages based on how well they served as source languages for the target language in other experiments Table 4).
For the tasks included in our experiments, choosing a source language based on the results for a similar task with similar input representations leads to only small losses: POS tagging results serve as a good predictor for parsing and vice versa, but not for topic classification. Similarly, the results for some topic classification settings are good predictors for each other, but not for POS tagging and parsing. These losses are often even smaller than when selecting source languages based on similarity measures (cf. Tables 3 and 4).
However, using other input representations weakens the predictive effect somewhat: e.g., for picking source languages for POS tagging with UDPipe, it is worse to choose based on de Vries et al.’s XLM-R results than to choose based on UDPipe’s parsing results. Topics-mbert is an especially poor predictor – this method likely suggests a fairly arbitrary mBERT language, which might or might not be similar to the target.
6 Conclusion
We investigated how linguistic and dataset level similarities impact cross-lingual transfer, and find that the most predictive similarity measures differ across tasks and input representations. For practitioners working on the tasks included in our study, we recommend choosing a source language based on a pertinent similarity measure, or, if results for an extensive transfer experiment involving a similar task and similar input representations are available, based on the transfer results in that experiment. Future work should investigate to what extent these patterns hold across even more languages, tasks, and NLP models.
Acknowledgements
We would like to thank Katherine Metcalf, Maureen de Seyssel, Sinead Williamson, Skyler Seto, Stéphane Aroca-Ouellette, and the anonymous reviewers for their helpful comments, discussions, and feedback.
Limitations
Throughout our study we have made a number of pragmatically motivated decisions. Although well motivated, these decisions come with some limitations that we address below.
Data
Despite the fact that our study includes more languages than any prior work, we were able to include only 263 of the world’s $\sim$7000 languages. Moreover, the usual high-resource languages are also over-represented in our work. We encourage data collection in more languages, especially those that are currently extremely low-resource or not available at all.
Despite our best efforts to mitigate any confounding effects, it is impossible to entirely avoid these. For instance, while the parallel nature of the SIB-200 datasets avoids some of the confounders related to UD (datasets containing texts from many different sources, genres, and domains), parallel datasets can also show translation artifacts Artetxe et al. (2020).
Models
We investigate one model type per task. Our model choices are motivated by practical considerations. We use UDPipe in part because of its state-of-the-art status for many languages and its wide coverage of languages, making it a likely out-of-the-box tool to be used for (morpho)syntactic annotations. An important motivation for us to choose MLPs for our topic classification experiments is the lightweight architecture that allows us to quickly train and compare many models, and we show that the performance is not far off from larger, more resource-intensive models (Table 2). We include $n$-gram-based MLPs in order to be able to have full control over the language model training data. Despite these models not having access to word order or syntactic information of the input data (and being more affected by exact word choices than a model with (sub)word embeddings would be), we include them because of their computational cheapness, and because there are no (comparable) monolingual embeddings for all $\sim$200 languages in the topic classification data.
We cannot guarantee that our findings hold for other models that could be used for the task, which is an avenue for investigation in future work. We also do not consider model- or tokenization-specific biases that may have different effects based on the word order White and Cotterell (2021), morphology Park et al. (2021), or writing system and orthography Sproat and Gutkin (2021).
Similarity measures
Within linguistics, there is a rich literature on defining similarity measures (cf. Borin, 2013). We chose to include similarity measures that are commonly used in NLP studies, and we adapt them where necessary. However, we acknowledge that there are many other similarity measures that are interesting to investigate.
Recent work has focused on extending lang2vec Khan et al. (2025); Amirzadeh et al. (2025) – we do not include these extensions in this paper as they were released concurrently to our work.
Tasks
We included POS tagging, dependency parsing, and topic classification in our study. An important motivation for this choice was the availability of data in a significant number of languages for these tasks. However, many more NLP tasks exist that are important to investigate. We echo Philippy et al. (2023): it is especially important to extend this work to more generative tasks, such as language modeling.
Analysis
The transfer results are overall rather symmetric (i.e., the scores when training on language A and evaluating on language B tend to be similar to when training on B and testing on A; compare the upper right and lower left triangles of the result heatmaps in Figure 2). However, these trends are not perfectly symmetrical. Transfer asymmetries have also been observed by other researchers Malkin et al. (2022); Protasov et al. (2024). However, we do not assume that these asymmetries mean that certain languages are intrinsically well-suited as source or target languages, but rather that they reflect peculiarities of the data sets (cf. Bjerva, 2024). We encourage further research on the (a)symmetries of cross-lingual transfer patterns.
Appendix A Languages and Resources
Resource Resource URL Universal Dependencies 2.14 Zeman et al. (2024) hdl.handle.net/11234/1-5502 SIB-200 Adelani et al. (2024) huggingface.co/datasets/Davlan/sib200 Grambank v1.0.3 Skirgård et al. (2023) zenodo.org/records/7844558 Glottolog v5.0 Hammarström et al. (2024) zenodo.org/records/10804582 ASJP v17 Wichmann et al. (2016) asjp.clld.org Jäger (2018a) osf.io/cufv7 lang2vec v1.1.6 (Littell et al., 2019) github.com/antonisa/lang2vec UDPipe 2 (2.12) (Straka, 2023) https://ufal.mff.cuni.cz/udpipe/2/models mBERT base uncased Devlin et al. (2019) huggingface.co/google-bert/bert-base-multilingual-uncased lme4 Bates et al. (2015) github.com/lme4/lme4 scikit-learn 1.5 Pedregosa et al. (2011) scikit-learn.org uroman 1.3.1.1 Hermjakob et al. (2018) github.com/isi-nlp/uroman GlotScript Kargaran et al. (2024) github.com/cisnlp/GlotScript SciPy Virtanen et al. (2020) scipy.org
Table 5: Details about the data, model, and software resources used in this paper.
Table 5 lists the versions of the datasets, models, and software we use. Lang2vec Littell et al. (2017) contains information on syntax, phonology, and phoneme inventories from WALS Dryer and Haspelmath (2013), SSWL Collins and Kayne (2011), Ethnologue Lewis et al. (2015), and PHOIBLE Moran et al. (2014). SIB-200 Adelani et al. (2024) is an annotated subset of FLORES-200 (NLLB Team et al., 2022), which in turn builds on previous versions and extensions of FLORES (Goyal et al., 2022; Guzmán et al., 2019; Doumbouya et al., 2023; AI4Bharat et al., 2023). UDPipe 2 (Straka, 2023) was trained on Universal Dependencies 2.12.121212http://hdl.handle.net/11234/1-5150 Its input representations include the last four layers of base-size uncased mBERT Devlin et al. (2019). This project uses the universal romanizer software ‘uroman’ written by Ulf Hermjakob, USC Information Sciences Institute (2015-2020).
[…]
Table 6: Languages and datasets used in our experiments. Continued and explained in next table.
[…]
Table 7: Languages and datasets used in our experiments, continued. ISO 639-3 codes in parentheses denote macrolanguage codes. Language family information is sourced from Glottolog Hammarström et al. (2024).
Tables 6 and 7 show the languages and datasets included in our experiments. Their geographic distribution is pictured in Figure 1.
A.1 Excluded UD treebanks
We exclude treebanks with code-switched language data (qaf_arabizi, qfn_fame, qtd_sagt, qte_tect), without text (ar_nyuad, en_gumreddit, gun_dooley, ja_bccwj, ja_bccwjluw), with glossed language (swl_sslc), and where the division into training and test split changed between releases 2.12 and 2.14 (tr_imst). We also exclude test sets with fewer than 20 sentences (kfm_aha, nap_rb, nyq_aha, soj_aha) and training sets with less than 100 sentences (bxr_bdt, hsb_ufal, kk_ktb, kmr_mg, lij_glt, olo_kkpp). We additionally exclude other training datasets that were not included in UDPipe 2 (ky_ktmu, de_lit, de_pud). Finally, we exclude the Czech UDPipe-2 models, as they use embeddings from a pretrained Czech model rather than mBERT.
A.2 Excluded SIB-200 languages
We exclude three Arabic dialects whose sentences in SIB-200 are almost identical to the Modern Standard Arabic (arb_Arab) sentences: ars_Arab, acm_Arab, acq_Arab. This is a known issue for FLORES-200, from which SIB-200 is derived.131313github.com/openlanguagedata/flores/issues/8
For the transliteration experiment, we exclude Japanese (jpn_Jpan) since uroman is not able to properly transliterate kanji as of version 1.3.1.1. We also exclude Mandarin with traditional characters (zho_Hant) and Cantonese (yue_Hant) since uroman crashes when trying to transliterate the corresponding datasets.
gb syn pho inv lex gen geo syn 0.61 pho 0.28 0.30 inv 0.22 0.16 0.30 lex 0.57 0.45 0.30 0.29 gen 0.56 0.47 0.31 0.25 0.87 geo 0.42 0.35 0.39 0.10 0.36 0.40 wor (UD) 0.44 0.35 0.24 0.39 0.59 0.47 0.17 swt (topics) 0.30 0.28 0.15 0.42 0.39 0.38 0.10 tri (topics) 0.16 0.24 0.12 0.37 0.24 0.24 -0.01 tri (top. tr.) 0.29 0.25 0.20 0.38 0.33 0.34 0.05 chr (UD) 0.26 0.38 0.35 0.28 0.29 0.28 0.23 chr (topics) 0.10 0.17 0.12 0.27 0.14 0.16 chr (top. tr.) 0.06 0.10 -0.03 0.16 0.04 0.06
Table 8: Correlations (Pearson’s r) between linguistic similarity measures (top) and between dataset similarity and linguistic similarity measures (bottom). Correlations with p-values >= 0.05 are replaced with —. Correlations between linguistic similarity measures are for the full set of languages included in our study. Key: gb=Grambank similarity, syn=syntactic similarity (lang2vec), pho=phonological similarity, inv=similarity of phoneme inventories, lex=lexical similarity, gen=phylogenetic relatedness, geo=geographic proximity, wor=word overlap, tri=character trigram overlap, chr=character overlap, top. tr.=transliterated topic classification data.
Appendix B Correlations between similarity measures
Table 8 shows how the different dataset-independent similarity measures are correlated with each other (top) and with the dataset-dependent similarity measures (bottom).
Appendix C Hyperparameters and Standard Deviations of Topic Classification Models
Parameter MLP ($n$-grams) MLP (mBERT) N-grams Min. length 1 Max. length 2, 3, 4 Type char, char_wb Max. features 5000, unlimited Max. epochs 5, 10, 20, 200, 300, 400 5, 10, 20, 30, 50, 200 Learning rate 0.0005, 0.001, 0.002 0.0005, 0.001, 0.002 Optimizer adam adam
Table 9: Hyperparameters used in the grid search for the topic classification models. Values in bold are the ones used in the final models.
Table 9 shows which hyperparameter values we included in the grid search for the topic classification models (the $n$-gram model trained and evaluated on the non-transliterated data, and the model using mBERT-based representations). We used the average development set scores for models (monolingually) evaluated on the following languages for hyperparameter tuning: arb_Arab, ayr_Latn, eng_Latn, eus_Latn, grn_Latn, kan_Knda, kat_Geor, kor_Hang, quy_Latn, vie_Latn, zho_Hans.
We calculate standard deviations for the final hyperparameter selection and the above-mentioned languages (evaluated monolingually on their development sets) over five random seeds. For the $n$-gram model, the standard deviations are between $0.0088$ (ayr_Latn) and $0.0286$ (kat_Geor), with a mean of $0.0181$. For the mBERT-based model, the standard deviations are between $0.0081$ (eng_Latn) and $0.0244$ (grn_Latn), with a mean of $0.0160$.
Appendix D NLP Task Results
D.1 Random baselines
For POS tagging and topic classification, we consider random baselines that randomly predict one of the 17 POS tags (or one of the seven topics) and are thus correct 5.9 % (or 14.3 %) of the time.
For each setup, all within-language performances (the diagonals in Figure 2) are above this threshold. Thus, all models learned something about the task, regardless of the training language and could be expected to be able to transfer some of that knowledge to a well-suited test language.
For some of the train–test language combinations, we see performances that are worse than random chance. This is the case for 3.1 % of the POS tagging results, 26.6 % of topics-base’s results, 14.5 % of topics-translit’s results, and 6.8 % of the results by topics-mbert. These worse-than-random transfer results are also meaningful: a POS tagger (or parser) trained on one language might learn grammatical patterns that run counter to how another language works, or a topic classification model might be mislead by superficial string overlaps between datasets that are false friends.
D.2 Heatmaps of transfer results
We include large, labelled versions of the heatmaps in Figure 2. We order the languages in each heatmap by clustering its rows (target languages) produced using Ward’s (1963) method, and applying the same order to the source languages. Because the heatmaps take up a lot of space, they are placed at the end of the appendix.
Figure 7 shows the POS tagging accuracies for all language pairs. Figures 8 (LAS) and 9 (UAS) present the parsing scores. Figure 10 shows the topic classification accuracies for topics-base, and Figure 11 for topics-translit. Finally, Figure 12 shows the results for the mBERT-based topic classification model.
D.3 Robustness across datasets of the same language
In UD, many languages have multiple training and/or test datasets: there are 124 training datasets in 70 languages and 268 test datasets in 153 languages. These datasets differ in the sentences that are annotated (oftentimes, they are from different sources and text genres) and often they are annotated by different people. Treebanks in the same language use the same writing system, with the exception of Sanskrit (and Mandarin Chinese if distinguishing between traditional and simplified characters). In SIB-200, there are eight languages that have two datasets each, which differ in their writing system.
We compare how robust the transfer patterns are across datasets of the same language. For test datasets, we calculate the correlation between the different models’ scores on a pair of datasets. For training datasets, we calculate the correlation between the evaluation scores produced by each pair of models trained on datasets of the same language.
Parsing
For parsing (LAS), most datasets of the same language produce very similar transfer patterns. Pearson’s $r$ and Spearman’s $\rho$ tend to produce very similar correlation numbers. Most correlations (both from the training and testing side) are above $0.9$, and most correlations below that are still above $0.8$. [^14] For the training datasets, we see exceptions ($r<0.8$) for two of the five Latin training sets (llct & proiel: $r=0.79$) and the two German training sets (hdt & gsd: $r=0.78$). For the test datasets, the following treebank pairs show correlations with $r<0.8$: two of the four Old East Slavic treebanks (birchbark & rnc: $r=0.79$), several of the seven Mandarin Chinese treebanks (hk & patentchar: $r=0.63$, beginner & patentchar: $r=0.63$, beginner & gsd: $r=0.69$. cfl & patentchar: $r=0.71$, gsdsimp & hk: $r=0.72$, beginner & gsdsimp: $r=0.72$, gsd & hk: $r=0.72$, cfl & gsd: $r=0.76$, beginner & pud: $r=0.79$, cfl & gsdsimp: $r=0.79$), the two Akkadian treebanks (pisandub & riao: $r=0.79$), the two Tamil treebanks (mwtt & ttb: $r=0.76$), and the two Sanskrit treebanks (ufal & vedic: $r=0.24$).
POS tagging
Most treebank pairs of the same language also show very similar transfer patterns. Again, most correlations are close to 1.151515For the training treebanks, we see correlations ($r$) below $0.8$ only for the following pairs: the two German training sets (hdt & gsd: $r=0.75$), the two Persian ones (perft & seraji: $r=0.73$), and the two Korean ones (gsd & kaist: $r=0.72$). For the test datasets this applies to some of the six Latin treebanks (circse & udante: $r=0.79$, circse & ittb: $r=0.79$), several of the eight Turkish datasets (atis & tourism: $r=0.55$, pud & tourism: $r=0.73$, penn & tourism: $r=0.74$, boun & tourism: $r=0.74$, gb & tourism: $r=0.78$, kenet & tourism: $r=0.79$), several of the seven Mandarin Chinese treebanks (hk & patentchar: $r=0.77$, beginner & patentchar: $r=0.78$, patentchar & pud: $r=0.79$, cfl & patentchar: $r=0.79$, gsd & patentchar: $r=0.79$), two of the four Old East Slavic treebanks (birchbark & rnc: $r=0.75$), the two Akkadian treebanks (pisandub & riao: $r=0.75$), and both Sanskrit treebanks (ufal & vedic: $r=0.57$).
topics-base
The writing system matters, and datasets in different writing systems (but the same language) show different transfer patterns. The two Mandarin Chinese datasets (traditional vs. simplified characters) show similar patterns (correlation as training sets: $r=0.96$, as test sets: $r=0.96$). For the test datasets, nearly all other correlations are insignificant, except for the Arabic and Devanagari Kashmiri datasets ($r=0.43$). For the training datasets, all correlations other than for the Mandarin datasets are either close to zero or negative. The strongest negative correlation is for Arabic- vs. Latin-script Modern Standard Arabic ($r=-0.58$).
topics-translit
For the transliterated version, all correlations are positive and/or close to zero, but not very high (the highest correlations are for the transliterated versions of the Tifinagh and Latin-script Tamasheq datasets: test $r=0.63$, train $r=0.34$). Note that we only have one transliterated version of the Mandarin Chinese datasets, as the transliteration tool did not work for the traditional script version. We hypothesize that the correlations for the transliterated data are fairly low since they do not necessarily have many $n$-grams in common. For instance, many of the languages with two datasets have one Arabic-script version and one Latin-script version. The latter contains vowels, while the automatically produced transliteration of the former only includes consonants.
topics-mbert
For topics-mbert, the correlations tend to be higher than for the $n$-gram-based models. For the training data, all correlations are above $0.6$ except for Achinese (which shows no significant correlation). For the test data, the results are more split, with mostly positive correlations, but also some negative ones (Minangkabau, $-0.24$; Banjar, $r=-0.32$). We hypothesize that the test data shows less coherent correlation patterns due to the inclusion of a language (in a given script) in mBERT’s test data having a stronger effect on the classification results than the inclusion of the training dataset’s language.
D.4 Effect of writing system
We compare transfer between languages using the same writing system to transfer across writing systems. For UD, we use GlotScript Kargaran et al. (2024) to determine the scripts; for SIB-200, writing system information is included as metadata.
Transfer between datasets with the same writing systems generally works better than between different scripts, however this is in part due to the language selection rather than the scripts themselves. Two-sample Kolmogorov–Smirnov tests indicate that the results within vs. across scripts come from different distributions (p-values all <0.0001; statistics: 0.51 for topics-base, 0.26 for topics-mbert, 0.16 for POS accuracy, 0.18 for LAS). However, the results of topics-translit for datasets that were in the same vs. different scripts before transliteration also come from different distributions (p <0.0001, statistic: 0.37), indicating that at least for SIB-200, the combinations of languages usually associated with the scripts alone already make an important difference.
Nonetheless, writing systems still play a role, especially for the $n$-gram-based models. Although topics-base and topics-translit achieve the same within-language accuracy (Table 2), topics-translit performs slightly better cross-lingually. Its within-dataset performance is slightly lower than for topics-base (69.4% vs. 70.3%), likely due to some language-specific information being lost when diacritics are removed. This would be compensated in the within-language performance by improved transfer between datasets of languages that originally had different scripts.
D.4.1 Effects of transliteration
Transliterated UD treebanks
Thirty-five UD test treebanks (in 25 languages) come with token-level Latin transliterations. We compare performance on the original data with performance on transliterated data. Figure 5 shows the performance differences of the POS taggers and parsers when evaluated on transliterated instead of original-script data. Performance is worse on most transliterated test treebanks, even when the models were trained on Latin-script treebanks. [^16] The exceptions are three languages whose scripts appear neither in any training treebanks nor in mBERT’s pretraining data (Amharic, Xibe, Sinhala), and one language that is usually written in Arabic, but unrelated to other languages using this writing system (Uyghur). This is in line with results from Pires et al. (2019), who observe that mBERT performs worse on transliterated than original-script data.
Topic classification
We compare the results of topics-base and topics-translit. Figure 6 shows the topic classification performance difference between the $n$-gram model trained and evaluated on the original data and the one trained and evaluated on transliterated data. For topics-translit, transfer between many original Cyrillic- and Latin-script language pairs is improved.
Eight of the languages in SIB-200 come in two script versions each. The writing systems are completely distinct, except for Mandarin Chinese, which has versions in traditional and simplified characters. For topics-base, transfer between the two scripts of a language is unsurprisingly always much lower than the performance on the same script (with accuracies between 8.3% and 25.9% for in-language cross-script transfer – excluding the Mandarin Chinese entries, for which cross-script accuracy is up to 64.7% – vs. within-language, within-script accuracies between 60.8% and 75.9% for the same languages). Although transfer between the transliterated versions of the datasets works better (between 14.2% and 50.5%), the accuracies are still much lower than the within-language, within-original-script accuracies (between 58.3% and 74.0%). This is likely due to different transliteration conventions for different writing systems (and due to missing vowels in the transliterated versions of abjads).
The patterns are also similar for topics-mbert, with within-language, cross-script scores between 8.3% and 45.6%, compared to within-language, within-script accuracies between 42.2% and 79.9%.
Figure 5: Differences between the performance on the original test treebanks and their transliterated counterparts for POS tagging (top) and parsing (LAS, bottom). Rows are for test sets, columns for training sets. Pink cells mark configurations where scores are better on the original data; green where scores are better on the transliterated treebank. Original writing systems are colour-coded. Writing systems in grey appear only for one language.
Figure 6: Differences between the n-gram-based topic classification performance on the original test languages and their transliterated counterparts. Rows are for test sets, columns for training sets. Pink cells mark configurations where scores are better on the original data; green where scores are better on the transliterated data. Original writing systems are colour-coded. Writing systems in grey appear only in one language.
Appendix E Correlations
E.1 Correlations between task results
POS LAS UAS topics base trans LAS 0.86 UAS 0.83 0.95 topics-base 0.39 0.43 0.40 topics-translit 0.40 0.53 0.48 0.68 topics-mbert 0.64 0.58 0.56 0.28 0.36
Table 10: Correlations between task results (Pearson’s r, all p-values are below 0.0001). Where possible, correlations are on a dataset level, otherwise on a language level.
Table 10 shows the correlations between task results. It is described in §4.2. The correlations across different task types only involve the 55 training and 84 test languages that appear both in UD and in SIB-200 (54 and 82 in the comparisons with the transliterated SIB-200 data).
E.2 Mixed effects models
[…]
Table 11: Linear mixed effects model results for each experiment. P-values are based on model comparison: *** = < 0.001, ** = < 0.01, * = < 0.05, . = < 0.1. Entries with p-values of >=0.05 are in grey. The last two rows show the number of training and test languages included in the analysis (i.e., the language pairs for which no fixed effects had missing information). “Corr” indicates pairs of strongly correlated fixed effects (see text, §E.2). “Same_script” is True iff the training and test datasets use the same writing system; “mbert_test” is True iff the test language is one of mBERT’s pretraining languages.
As described at the end of §5.1, we fit one linear mixed effects model per experiment. We model the NLP results (POS accuracy, parsing scores, topic classification accuracy) as the dependent variable, and the training and test languages as random effects. The fixed effects are the similarity measures, as well as binary variables (dummy-coded) indicating whether the training and test datasets have the same writing system and whether the test language is among mBERT’s pretraining languages. Because the models can only be fit for entries where no data points are missing (i.e., geo, lex, syn, pho, inv, and gb are all defined for the language pair at hand), the number of language pairs included in each mixed effects analysis is much smaller than for the correlations calculated independently per effect in §5.1.
Collinearity was observed between several fixed effects in all models (with correlation coefficients between $-0.769$ and $-0.819$ for phylogenetic relatedness and lexical similarity, between $-0.442$ and $-0.496$ for character overlap and sharing the same writing system, and between $-0.447$ and $-0.509$ for gb and syn). We mark them as such in Table 11, which shows the estimates and their significance values. We report the significance values based on model comparison (i.e., by comparing the full model and a model with one predictor taken out) and thus significance values are robust to collinearity.
The values for these related effects should be interpreted with these correlations in mind, as this can impact the estimates for these variables.
E.3 Correlations between NLP results and similarity measures
Tables 12 and 13 show the correlations between POS/parsing scores and the similarity measures for each test language. Tables 14, 15, and 16 show the same for the topic classification experiments. Because these tables take up a lot of space, they are placed at the very end of the appendix.
Figure 7: POS tagging accuracy scores for all combinations of source (columns) and target languages (rows), ordered by target language clusters (Ward’s method). The darker a cell, the better the score.
Figure 8: Labelled attachment scores for all combinations of source (columns) and target languages (rows), ordered by target language clusters (Ward’s method). The darker a cell, the better the score.
Figure 9: Unlabelled attachment scores for all combinations of source (columns) and target languages (rows), ordered by target language clusters (Ward’s method). The darker a cell, the better the score.
Figure 10: Topic classification accuracy scores (MLP with n-grams, original writing systems) for all combinations of source (columns) and target languages (rows), ordered by target language clusters (Ward’s method). The darker a cell, the better the score.
Figure 11: Topic classification accuracy scores (MLP with n-grams, transliterated data) for all combinations of source (columns) and target languages (rows), ordered by target language clusters (Ward’s method). The darker a cell, the better the score.
Figure 12: Topic classification accuracy scores (MLP with mBERT representations) for all combinations of source (columns) and target languages (rows), ordered by target language clusters (Ward’s method). The darker a cell, the better the score.
Test language POS LAS lang script mBERT size pho inv geo syn gb gen lex chr wor size pho inv geo syn gb gen lex chr wor abk Cyrl -36* N/A 29 30 -30 N/A 54* 30* abq Cyrl N/A N/A N/A N/A 24 N/A N/A N/A N/A afr Latn x N/A 31* N/A N/A 52* 59* 46* 36* N/A 45* 37* N/A N/A 71* 78* 46* 56* aii Syrc -32* N/A N/A N/A N/A 29 -23 N/A N/A N/A N/A ajp Arab -32* N/A N/A 30 N/A [36] N/A N/A 30 N/A [51*] akk Latn -25 N/A N/A N/A -25 N/A N/A N/A 24 aln Latn -37* N/A N/A N/A 32* 23 N/A N/A 40* N/A 54* 31* als Latn 32* 50* 46* 25 48* 29 41* 81* 66* 27 37* 41* amh Ethi -31* apu Latn -40* 35* -33* 28 23 aqz Latn -40* N/A N/A -37* N/A N/A arb Arab x N/A N/A 27 53* 42* 29 28 N/A N/A 77* 53* 25 35* 33* arr Latn -33* N/A 36 N/A 27 N/A 40* -60* N/A 30 aze Latn x -31* N/A N/A 34* 31* 61* N/A 35* N/A azz Latn -39* N/A N/A 31* N/A 37* 37* -27 N/A N/A 45* N/A 49* 37* bam Latn -40* 45 37* 30 -37* 26 39* bar Latn x N/A N/A 24 N/A N/A 48* N/A 37* 37* N/A N/A 40* N/A N/A 58* N/A 49* 42* bej Latn -34* 25 54* bel Cyrl x N/A N/A 37* N/A 64* 58* 56* 50* 34* N/A N/A 51* N/A 66* 65* 66* 63* 38* ben Beng x -24 N/A 26 39* N/A 35* 38* 32* bho Deva -40* N/A N/A 25 47* 29 33* 32* 33* N/A N/A 48* 81* 57* 43* 58* 54* 60* bor Latn -42* 40 N/A 29 36* -32* N/A 30 bre Latn x -31* 37* 51* 38* 27 53* 46* 31 48* 76* 64* 39* 52* 36* bul Cyrl 39* 36* 71* N/A 51* 47* 40* 25 47* 48* 84* N/A 54* 54* 47* 29 bxr Cyrl -35* N/A N/A 28 N/A 34 -24 N/A N/A 34* N/A 70* cat Latn x 44* 72* 61* 64* 67* 53* 48* 41* 52* 83* 77* 70* 74* 51* 56* ceb Latn x -35* N/A -34* N/A N/A 38* 27 N/A -48* N/A N/A 32* ces Latn x N/A 40* 36 30 35* 33* 52* 52* N/A 39* 51* 54* 50* 51* 54* 51* 56* chu Cyrl N/A N/A 26 N/A N/A 49* 46* 62* 63* N/A N/A 38* N/A N/A 60* 62* 73* 79* ckt Cyrl -32* -38* -33 -29 cmn Hans, Hant x 44 54* 51* 40* 29 29 62* 33 64* 52* 38* 55* 32* 39* cop Copt N/A N/A 35 69* 87* 91* 90* 92* N/A N/A 25 46* 73* 93* 98* 98* 99* cpg Grek -38* N/A N/A 26 N/A N/A 36* N/A 42* 43* N/A N/A 43* N/A N/A 53* N/A 59* 43* cym Latn x -30 N/A N/A 34* 42* 41* 39* 54* 50* 48* N/A N/A 49* 71* 73* 51* 66* 53* 58* dan Latn x N/A N/A 33* 49* 59* 45* 56* 57* 48* N/A N/A 45* 71* 77* 55* 63* 57* 55* deu Latn x 34 50* N/A 50* 55* 35* 50* 50* 50* 36* 70* N/A 60* 65* 48* 62* egy Latn N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A 41* 27 ekk Latn x N/A N/A 25 N/A 32 36* 37* 58* 43* N/A N/A 42* N/A 52* 53* 55* 56* 61* ell Grek x 47 32 34* 62* 63* 43* 37* 45* 24 48 42* 49* 82* 76* 48* 39* 55* 27 eme Latn -39* N/A 23 N/A 33* 37* -34* N/A N/A eng Latn x 44 33 29 37 39* 38* 33* 42* 50* 42* 32* 57* 64* 52* 51* 42* 52* ess Latn 36 N/A N/A -28 N/A N/A 35* eus Latn x 53* 44* N/A 41* 52* 43* 42* 75* 72* N/A 71* 43* 73* fao Latn N/A N/A 45* N/A 52* 52* 58* 61* 37* N/A N/A 46* N/A 66* 62* 68* 53* 39* fin Latn x 42 25 37 34* 36* 59* 41* 47 31 39* 61* 36 44* 46* 56* 52* fra Latn x 37* 51* 50* 63* 59* 42* 39* 34 46* 69* 69* 70* 67* 42* 46* frm Latn -23 N/A N/A 33* N/A N/A 63* N/A 26 35* N/A N/A 44* N/A N/A 72* N/A 28 40* fro Latn -24 N/A N/A 36* N/A N/A 66* N/A 24 45* N/A N/A 42* N/A N/A 72* N/A 46* gla Latn -32* N/A N/A 26 41* N/A 54* 74* 41* 68* N/A N/A 41* 64* N/A 67* 85* 46* 86* gle Latn x -30 N/A 50* 38* 50* 51* 43* 55* 51* 57* N/A 56* 52* 76* 73* 57* 62* 50* 65* glg Latn x N/A 43* N/A 56* 62* 66* 45* 46* N/A 37 53* N/A 75* 70* 72* 47* 51* glv Latn -34* N/A N/A 28 N/A N/A 55* 75* 34* 68* N/A N/A 39* N/A N/A 71* 84* 36* 81* got Latn -25 N/A N/A N/A N/A 29 42* 45* 74* N/A N/A 24 N/A N/A 49* 62* 47* 96* grc Grek -28 N/A N/A 27 N/A 45* 45* 57* 51* 48* N/A N/A 48* N/A 68* 64* 71* 66* 62* grn Latn -33* 30 47* N/A N/A 40* 37* 41 41* N/A N/A gsw Latn N/A N/A N/A N/A 45* 49* 32* N/A N/A 25 N/A N/A 56* 62* 36* gub Latn -41* N/A N/A N/A N/A 38* N/A N/A 43* guj Gujr x N/A N/A N/A 31* 29 29 N/A 28 N/A N/A 40* 40* 28 gun Latn -43* N/A N/A 27 -24 N/A N/A hat Latn x -26 N/A N/A 29 N/A 31* N/A 48* 37* N/A N/A 38* N/A 38* 39* N/A 48* 27 hau Latn -43* 50* N/A -39* 43 N/A 26 hbo Hebr -37* N/A N/A 36* N/A 50* 46* 49* 41* 33* N/A N/A 49* N/A 70* 55* 56* 55* 46* heb Hebr x N/A 26 61* 41* 27 47* 33* N/A 35 33* 83* 54* 23 32* 53* 37* hin Deva x 51* 34 67* 50* 41* 43* 41* 33* 59* 36* 74* 73* 63* 64* 53* 58* hit Latn -24 N/A N/A N/A N/A -28 N/A N/A N/A N/A hrv Latn x N/A 39* N/A N/A 37* 31 60* 46* N/A 44* 49* N/A N/A 54* 52* 55* 56* hsb Latn -27 N/A N/A 36* N/A N/A 31* 29 50* 53* N/A N/A 48* N/A N/A 49* 52* 58* 51* hun Latn x 42 25 31 33 28 31 58* 40* 49 40* 59* 53* 42* 47* 53* 53* hye Armn x 45 60* 51* 38* 43* 51* 34* 58* 43* 36* 58* 56* 55* 61* 68* 51* hyw Armn N/A N/A N/A N/A N/A 42* 46* 55* 38* N/A N/A N/A N/A N/A 58* 63* 70* 55* ind Latn x 46 41* N/A 31* 29 58* 42* 40 62* N/A 39* 43* 52* 49* isl Latn x N/A 30 41* 46* 45* 54* 59* 60* 40* N/A 42* 46* 66* 59* 61* 64* 56* 45* ita Latn x N/A 35* 59* 58* 58* 61* 44* 42* N/A 36 44* 79* 73* 64* 67* 45* 48* jaa Latn -36* N/A 28 N/A 27 23 -37* N/A N/A jav Latn x -24 N/A N/A 25 58* 44* -26 N/A N/A 29 25 54* 47* jpn Hira x -24 53* 32* 38* 41* 48* 48* 41* 46* 65* 45* 51* 65* 77* 81* 77* 71* 78* kat Geor x 73* 34 -28 30* 27 71* 31 -31 37* kaz Cyrl x -27 N/A N/A N/A N/A 36* N/A N/A 25 N/A N/A kir Cyrl x -27 N/A N/A 30* 31* N/A N/A kmr Latn -43* N/A 29 32* -39* 41 N/A 24 24 koi Cyrl -39* N/A N/A 27 N/A 35* -25 N/A N/A 38* N/A 30 25
Table 12: Correlations between POS/LAS results and similarity measures. Continued/explained in next table.
Test language POS LAS lang script mBERT size pho inv geo syn gb gen lex chr wor size pho inv geo syn gb gen lex chr wor kor Hang x 32* 55* 52* 35* 40* 34* 37* 46 31 36* 79* 73* 57* 59* 58* 58* kpv Cyrl -38* 24 N/A 36* -28 42 36* N/A 31* 37* krl Latn -27 N/A N/A N/A N/A 34* 36* 53* 50* N/A N/A 35* N/A N/A 41* 41* 54* 41* lat Latn x -28 N/A N/A 26 N/A 41* 51* 33* 41* N/A N/A 44* N/A 43* 56* 65* 37* 58* lij Latn -29 N/A N/A 42* N/A N/A 60* 63* 43* 44* N/A N/A 52* N/A N/A 64* 66* 45* 41* lit Latn x -24 50* 28 35 36* 57* 32* 40 47* 48* 47* 52* 42* 52* 56* 40* ltg Latn -23 N/A N/A 28 N/A N/A N/A 46* 23 N/A N/A 41* N/A N/A 32* N/A 43* ltz Latn x N/A N/A N/A N/A 39* 50* N/A N/A N/A N/A 46* 58* 28 lvs Latn x N/A N/A 28 N/A [—] 27 35* 55* 37* N/A N/A 46* N/A [36] 44* 52* 56* 46* lzh Hant 26 N/A N/A 47* N/A N/A 65* N/A 57* 65* 47* N/A N/A 44* N/A N/A 85* N/A 60* 87* mal Mlym x N/A N/A 35* N/A N/A 48* 23 28 35* mar Deva x N/A 64* 53* 29 36* 33* 27 N/A 32* 69* 65* 43* 49* 47* 42* mdf Cyrl -35* N/A N/A 29 N/A 30 33* 40* 35* 44* N/A N/A 43* N/A 44* 59* 62* 62* 71* mkd Cyrl x N/A 36 37* N/A 69* 49* 48* N/A 46* 51* N/A 72* 50* 50* mlt Latn -32* N/A 34 N/A 31 47* 60* 27 76* N/A 49* 27 N/A 59* 69* 82* 29 97* mpu Latn -36* N/A N/A N/A N/A N/A N/A myu Latn -39* N/A N/A 32* -39* N/A N/A 31* myv Cyrl -33* N/A N/A 29 N/A 45* 51* 61* 46* 57* N/A N/A 38* N/A 60* 81* 85* 71* 88* nds Latn x N/A 28 N/A N/A 40* 53* 39* 48* N/A 39* N/A N/A 50* 63* 42* 47* nhi Latn -44* N/A N/A 24 N/A 29 41* -37* N/A N/A 29 N/A 30 42* nld Latn x N/A 36 31* 56* 59* 48* 58* 53* 48* N/A 49* 41* 77* 76* 61* 68* 55* 56* nno Latn x N/A N/A 32* N/A N/A 52* 62* 59* 56* N/A N/A 43* N/A N/A 61* 68* 57* 64* nob Latn x 32* 49* N/A 50* 59* 60* 55* 42* 73* N/A 58* 65* 59* 62* olo Latn -36* N/A N/A N/A 34* N/A 46* 41* N/A N/A 35* N/A 35 48* N/A 46* orv Cyrl N/A N/A 30 N/A N/A 59* N/A 44* 51* N/A N/A 47* N/A N/A 72* N/A 53* 59* ota Latn -35* N/A N/A N/A N/A 34* N/A 32* 42* N/A N/A N/A N/A 57* N/A 54* otk Orkh N/A N/A N/A N/A N/A N/A 68* 32* N/A N/A N/A N/A N/A N/A 29 pad Latn -41* N/A 37* -39* N/A pcm Latn N/A 38* N/A N/A 63* N/A 31* 77* N/A 37 40* N/A N/A 67* N/A 37* 91* pes Arab x 48 25 39* 44* 40* 26 28 67* 45* 53* 46* 57* 55* 43* 45* pol Latn x 44 37* 44* 48* 34* 29 54* 39* 34 49* 69* 62* 49* 47* 52* 45* por Latn x N/A 31 43* 53* 51* 59* 59* 51* 50* N/A 42* 52* 78* 71* 66* 64* 52* 54* qpm Latn -33* [54*] [—] [35*] [40*] N/A 42* [27] 52* 51* [47] [43*] [49*] [68*] N/A 59* [48*] 50* 59* quc Latn -42* N/A N/A 25 N/A 32 52* N/A 42* 27 ron Latn x -26 45 32* 44* N/A 55* 59* 50* 48* 33 44* 76* N/A 63* 65* 49* 53* rus Cyrl x 42* 61* 52* 56* 54* 43* 35* 40 53* 77* 54* 65* 64* 55* 40* sah Cyrl -36* 50* 37 32* 27 25 -28 24 81* 54* 40* 38* san Deva, Latn -41* N/A 47* 25 N/A N/A 30* 39* -25 N/A 62* 56* N/A N/A 27 56* 48* 66* say Latn -42* N/A N/A -39* N/A N/A 30 sga Latn -38* N/A N/A N/A N/A 31* 45* 27 37* -34* N/A N/A 47* N/A N/A 48* 60* 33* 39* sin Sinh -25 32* N/A N/A -46* 33 42* N/A N/A sjo Mong N/A N/A 34* N/A N/A -35* N/A N/A 34* N/A N/A slk Latn x N/A N/A 39* N/A N/A 32* 28 65* 41* N/A N/A 49* N/A N/A 51* 51* 59* 48* slv Latn x N/A 39* 39* 47* 40* 43* 40* 48* 52* N/A 49* 49* 69* 56* 53* 53* 47* 58* sme Latn -34* N/A N/A 25 N/A 40* 64* 66* 44* 62* N/A N/A 28 N/A 58* 91* 93* 41* 96* sms Latn -43* N/A N/A 24 N/A 30 36* 34* 34* 39* -27 N/A N/A 38* N/A 47* 37* 35* spa Latn x 43* 60* N/A 63* 64* 44* 52* 39 51* 78* N/A 70* 70* 44* 56* srp Latn x N/A N/A 35* N/A N/A 36* N/A 61* 43* N/A N/A 47* N/A N/A 53* N/A 55* 53* swe Latn x N/A 43* 27 59* 65* 48* 57* 57* 43* N/A 47* 39* 78* 79* 57* 62* 57* 48* tam Taml x N/A 23 50* 48* 33* 36* 38* 31* N/A 37* 70* 65* 39* 41* 40* 37* tat Cyrl x -26 N/A N/A 40 36 N/A N/A 65* 55* tel Telu x 50* 43* 29 34* 31* 28 35 39* 81* 65* 38* 43* 36* 35* tgl Latn x -32* 48 -37* 33 37* 25 -46* 74* 31* tha Thai 44* -52* -29 -30* 60* 30* 28 tpn Latn -39* N/A N/A 23 29 -36* N/A N/A 26 tsn Latn -33* N/A N/A N/A 36* 31* -27 N/A N/A N/A 30 33* tur Latn x -28 36 38* 34* 31 43* 42* 76* 71* 55* 53* 52* uig Arab -35* N/A N/A 34* 63* N/A 60* 61* 38* 58* N/A N/A 42* 76* N/A 79* 78* 61* 77* ukr Cyrl x N/A 41* 37* 68* 52* 57* 54* 48* 30 N/A 51* 51* 82* 60* 65* 65* 61* 36* urb Latn -27 N/A 29 -30 53* N/A 26 urd Arab x N/A 43* 29 71* N/A 47* 51* 28 37* N/A 62* 48* 74* N/A 70* 77* 39* 64* vep Latn -35* N/A N/A N/A 40* 44* 44* 41* N/A N/A 41* N/A 34 61* 64* 44* vie Latn x -31* 40 39* 31 35* 34* 54* 37* 51* 50* 59* 32 42* 39* 58* 43* wbp Latn -38* N/A N/A 29 N/A N/A 34* 70* 32* wol Latn -31* 43 55* 35* 34 71* 72* 76* 43* 73* 65* 43* 50* 82* 96* 95* 36* 96* xav Latn -46* N/A N/A 28 -36* N/A N/A xcl Armn -36* N/A N/A 31* N/A 32* 46* 39* 37* N/A N/A 42* N/A 42* 43* 39* 37* xnr Deva -36* N/A 28 N/A N/A N/A 38* 38* N/A 41* N/A N/A 30 N/A 50* 47* xum Latn -28 N/A N/A N/A N/A N/A 32* 28 -25 N/A N/A N/A N/A N/A 24 yor Latn x -36* 40 N/A 47* 26 61* N/A 47* yrl Latn -42* N/A N/A 31* 42* -34* N/A 38* N/A 47* 44* yue Hant 36 30 36* 53* 28 mean (all) -19 17 9 17 37 28 26 29 32 33 -7 17 20 27 57 42 35 38 36 34 -21;-15 11;23 5;12 13;19 31;42 22;32 22;29 24;33 28;35 30;36 -8;-4 10;22 15;25 23;30 50;63 35;47 30;40 32;43 32;39 29;38 mean (mBERT) -8 23 11 20 46 40 34 38 43 35 0 24 28 31 69 60 46 49 44 40 -11;-4 14;30 6;15 15;24 41;51 34;46 28;38 32;42 38;46 31;38 0;0 14;32 22;34 25;36 64;72 55;64 40;50 43;54 40;48 35;45 mean ($\neg$¬mBERT) -27 8 7 14 20 14 20 21 23 32 -12 6 9 24 36 23 27 27 30 29 -30;-23 0;15 2;12 10;17 10;29 7;20 14;25 14;27 17;28 26;36 -15;-8 0;12 2;15 19;28 20;50 13;31 20;34 19;35 24;34 21;35
Table 13: Correlations (Pearson’s r $\times$ 100, to save space) between POS/LAS results and similarity measures for each test language, continued. Where a training or test language has multiple datasets, we use language-wise score averages for calculating the correlations. The asterisk* denotes correlations with a p-value below 0.01. Grey cells with a bar (—) denote correlations with a p-value of 0.05 or above. Square brackets [ ] mean that no entry for this ISO code was found in the linguistic databases, so the entry for its macrolanguage code was used instead (bul for qpm, apc for ajp, lav for lvs). ‘N/A’ means that no correlation score could be calculated due to missing entries in the linguistic databases. The bottom rows show the mean scores across test languages (‘—’ entries are treated as zeros, ‘N/A’ entries are ignored). Below each row of mean scores are the corresponding 95% confidence intervals (numbers separated by semicolons). ‘avg mBERT’ (‘avg $\neg$¬mBERT’) is for the test languages that are (are not) in mBERT’s pretraining data.
Test language Original Transliterated mBERT lang script   mB. pho inv geo syn gb gen lex chr tri pho inv geo syn gb gen lex chr tri pho inv geo syn gb gen lex chr swt ace Arab, Latn N/A 33* 38* 52* 49* 41* 60* N/A 26* 28* 44* 56* 53* 40* N/A -25* 33* 53* 28* aeb Arab N/A N/A N/A N/A 60* 68* 55* 81* N/A N/A 16 N/A N/A 75* 80* 32* 90* N/A N/A 24* N/A N/A 20* afr Latn x N/A 36* N/A N/A 43* 51* 55* 69* N/A 34* -14 N/A N/A 48* 54* 14 54* N/A -33* N/A N/A 39* 31* ajp Arab N/A N/A 17 N/A [37*] 61* 66* 49* 78* N/A N/A 29* N/A [57*] 79* 82* 35* 89* N/A N/A 25* N/A [20] 18* aka Latn 44* 22* 44* N/A 53* 45* 61* 72* 36* 22* 38* N/A 57* 59* 23* 46* 34* 33* N/A 40* 14 52* 58* als Latn 29 39* 26* 18 22* 34* 62* 76* 42* 43* 28* 39* 41* 36* 43* 56* 44* 32* 48* 47* 46* 28* amh Ethi 28 29* 47* 39* 49* 49* 55* 29 22 22* 40* 55* 40* 50* 33* 15 25* 40* apc Arab N/A N/A 19* 38* 61* 66* 55* 80* N/A N/A 28* 32* 58* 78* 82* 41* 89* N/A N/A 27* 20 17 arb Arab, Latn x N/A N/A 16 26* 56* 59* 22* 43* N/A N/A 31* 27 53* 77* 80* 19* 81* N/A N/A 14 27 31* 20* ary Arab N/A 23* 24 N/A 59* 66* 54* 79* N/A 37* 40* N/A 77* 81* 32* 89* N/A 22* 29 N/A 15 22* arz Arab 30 26* 18 24 35* 61* 70* 53* 81* 37* 38* 26* 35* 64* 78* 84* 21* 90* 35* 23* 23 19* asm Beng N/A 26* 34 N/A 36* 52* 36* 61* N/A 50* 52* N/A 66* 76* 63* N/A N/A 31* 21* 19* ast Latn x N/A N/A 32* N/A N/A 65* 68* 56* 73* N/A N/A 43* N/A N/A 70* 73* 16 64* N/A N/A 31* N/A N/A 43* 37* awa Deva N/A 34* 42* N/A 49* 64* N/A 77* 86* N/A 51* 51* N/A 56* 72* N/A 23* 75* N/A N/A 40* 33* N/A 25* 17 ayr Latn 39* 39* 34* 53* 49* 52* 64* 29 30* 25* 46* 61* 57* 38* 30* 40* 20 14 17 46* 46* azb Arab x 22* 48* N/A 16 22* 54* 71* 34* 32* N/A 42* 50* 41* 78* 36* 23* N/A 14 18* azj Latn N/A 35* N/A 28* 37* 32* 53* 68* N/A 56* 31* N/A 42* 63* 67* 17 54* N/A 21 34* N/A 21 15 bak Cyrl x 39* 30* 30* 42* 31* 52* 49* 60* 80* 52* 34* 37* 61* 53* 64* 65* 26* 57* 48* 41* 41* 25* 16 18 23* 14 bam Latn 27 41* 17 80* 35* 47* 51* 54* 64* 29* 17 87* 30* 61* 59* 39* 40* 34* 16 24* 55* 60* ban Latn N/A 45* N/A 46* 59* 55* 60* 74* N/A 42* 22* N/A 43* 60* 58* 56* N/A -18 N/A -19 17 bel Cyrl x N/A N/A 22* N/A 35* 36* 41* 52* 71* N/A N/A 23* N/A 42* 54* 63* 26* 63* N/A N/A 41* N/A 55* 44* 35* 23* bem Latn N/A 45* 29* N/A 51* 55* 58* 52* 63* N/A 33* 31* N/A 57* 55* 61* 47* N/A 39* 38* N/A 54* 59* 53* 50* 62* ben Beng x 34* 27* 81* N/A 37* 49* 39* 64* 31* 56* 53* 77* N/A 69* 78* 36* 63* N/A 29* 22* bho Deva N/A N/A 40* 45* 54* 65* 58* 67* 83* N/A N/A 50* 45* 59* 72* 73* 71* N/A N/A 45* 30* 34* 26* 36* 22* bjn Arab, Latn N/A N/A 24* N/A N/A 61* 63* 35* 55* N/A N/A 42* N/A N/A 69* 67* 56* N/A N/A -18* N/A N/A 43* 15 bod Tibt 24 23* 43 34* 62* 56* 34* 65* 32* 17 60* 47* 76* 71* 72* 58* 33* 35* 36* 20* 48* bos Latn x N/A N/A 21* 29 N/A 31* 40* 64* 79* N/A N/A 36* 61* N/A 66* 73* 28* 72* N/A N/A 36* 63* N/A 44* 35* 14 bug Latn N/A 47* N/A 46* 52* 49* 64* 77* N/A 40* N/A 44* 55* 53* 14 55* N/A -19* N/A 15 16 47* 38* bul Cyrl 33* 19* 22 N/A 39* 40* 51* 75* 30 59* 41* 51* N/A 61* 67* 34* 66* 27 39* 49* N/A 46* 38* 24* 16 cat Latn x 48* 29* 49* 44* 60* 62* 59* 72* 24 43* 42* 57* 58* 69* 70* 65* 27 33* 34 49* 43* 39* ceb Latn x N/A 44* N/A N/A 62* 62* 60* 74* N/A 42* 24* N/A N/A 66* 68* 60* N/A N/A N/A 14 20* ces Latn x N/A 40* 18* 24 29* 37* 67* 82* N/A 50* 37* 57* 43* 59* 66* 20* 72* N/A 41* 52* 54* 46* 36* 17 cjk Latn N/A N/A 29* N/A N/A 50* 57* 49* 61* N/A N/A 24* N/A N/A 44* 55* 46* N/A N/A 36* N/A N/A 51* 43* 59* 64* ckb Arab N/A N/A 18* N/A N/A 31* 50* 30* 60* N/A N/A 27* N/A N/A 43* 54* 15 54* N/A N/A N/A N/A 19* 18 29* cmn Hans, Hant x 27* 15 30* 30* 56* 56* 69* 24 41* 17 28* 67* 68* 27* 25 26* 15 36* 25* 30* 34* 42* 37* crh Latn N/A N/A N/A 29* 24* 28* 59* 70* N/A N/A 29* N/A 49* 66* 70* 25* 54* N/A N/A 33* N/A 15 19* cym Latn x N/A N/A 24 25* 28* 45* 45* 57* N/A N/A 22* 22 36* 35* 51* 19* 40* N/A N/A 35* 41* 37* 31* 21* dan Latn x N/A N/A 26* 38* 37* 39* 51* 59* 73* N/A N/A 38* 58* 50* 48* 56* 21* 62* N/A N/A 43* 61* 54* 44* 33* 14 deu Latn x 30* 36* 24* 35* N/A 45* 52* 53* 66* 52* 45* 37* 56* N/A 48* 53* 15 55* 47* 41* 57* N/A 41* 32* dik Latn N/A N/A N/A N/A 40* 45* 49* 62* N/A N/A N/A N/A 49* 50* 26* N/A N/A N/A N/A 18 53* 48* dyu Latn N/A 43* N/A 34* 40* 49* 54* 64* N/A 31* 15 N/A 36* 54* 57* 49* N/A 24* 23* N/A 20* 26* 58* 55* dzo Tibt N/A N/A 21* N/A 40* 64* 60* 63* 66* N/A N/A 14 N/A 38* 78* 75* 56* 72* N/A N/A N/A 30* 29* 32* 30* 32* ekk Latn x N/A N/A N/A 37* 38* 62* 74* N/A N/A 29* N/A 38* 43* 45* 17 52* N/A N/A 45* N/A 44* 16 ell Grek x 19 21 29* 43* 15 47* 31* 41* 18 43* 42* 46* 57* 21* 54* 36* 34* 52* 49* 44* 29* 18 eng Latn x 23 32* 25* 19 35* 35* 47* 61* 34* 17 30* 31* 38* 35* 49* 35* 41* 45* 45* 34* 31* epo Latn N/A N/A 28* N/A N/A 22* N/A 60* 74* N/A N/A 38* N/A N/A 25* N/A 16 64* N/A N/A 28* N/A N/A N/A 30* 23* eus Latn x 36* 18* 27* N/A 28* 62* 77* 45* 29* 21 40* N/A 36* 59* 32* N/A ewe Latn 33* 51* 20* 43* 50* 48* N/A 59* 70* 25 44* 21* 33* 51* 50* N/A 17 51* 40* 29* 24 21 43* N/A 55* 61* fao Latn N/A N/A 22* N/A 39* 45* 52* 53* 66* N/A N/A 30* N/A 48* 56* 61* 17 51* N/A N/A 34* N/A 37* 36* 27* 24* fij Latn 25 41* 17 32* 49* 44* 45* 47* 56* 27* 21* 26 45* 45* 48* 20* 26* 28* 25 23* 22* 56* 55* fin Latn x 49* 22 27* 42* 40* 55* 69* 33* 51* 24* 48* 41* 49* 48* 19* 49* 33* 48* 59* 43* 17 fon Latn N/A 40* 20* 53 38* 53* 54* 43* 55* N/A 25* 16 52 33* 48* 64* 34* N/A 31* 32* 25* 43* 23* 44* 51* fra Latn x 41* 26* 39* 41* 53* 53* 58* 71* 46* 34* 35* 48* 47* 57* 57* 59* 46* 40* 47* 47* 40* 35* fur Latn N/A 40* 28* N/A N/A 60* 62* 55* 68* N/A 40* 37* N/A N/A 66* 66* 59* N/A 33* N/A N/A 40* 37* 25* 20* fuv Latn N/A 44* N/A 41* 48* 43* 51* 60* N/A 32* N/A 40* 51* 55* 36* N/A 24* 20* N/A 45* 32* 21* 58* 56* gaz Latn N/A N/A N/A N/A 47* 56* 24* 34* N/A N/A 20* N/A N/A 51* 64* 31* N/A N/A N/A N/A 49* 53* gla Latn N/A N/A N/A 40* 66* 28* N/A N/A 25 N/A 52* 71* 26* N/A N/A 26* 28* N/A 29* 24* 47* 27* gle Latn x N/A 50* 29* 27* 39* 62* 20* 37* N/A 40* 17 33* 42* 54* 70* 30* N/A 33* 42* 37* 25* 26* glg Latn x N/A 43* 33* N/A 49* 62* 64* 57* 73* N/A 34* 40* N/A 59* 68* 67* 14 66* N/A 31* N/A 49* 41* 35* grn Latn 34* 51* 37* 59* N/A 35* N/A 61* 75* 33* 39* 28* 61* N/A 43* N/A 62* 30* 25* 32* 41* N/A 15 N/A 52* 46* guj Gujr x N/A 20 N/A N/A 21* 34* 21* 48* N/A 54* 52* N/A N/A 65* 71* 36* 67* N/A 15 N/A N/A 33* 26* hat Latn x N/A N/A 33* N/A 35* 32* N/A 54* 65* N/A N/A 32* N/A 22 27* N/A 18* 45* N/A N/A 18 N/A 36* N/A 24* hau Latn 47* 14 27* N/A 38* 46* 43* 53* 39* 24 N/A 49* 56* 14 25* 25* 18* N/A 53* 52* heb Hebr x N/A 21 21 33* 42* 51* N/A 22* 28* 35* 50* 59* 19* 68* N/A 24* 25 21 hin Deva x 34* 40* 44* 44* 62* 56* 74* 85* 33* 59* 50* 49* 47* 70* 72* 30* 74* 25 41* 38* 32* 30* 22* 16 hne Deva N/A N/A 40* N/A N/A 63* N/A 78* 86* N/A N/A 55* N/A N/A 73* N/A 29* 73* N/A N/A N/A N/A 32* N/A 29* 20* hrv Latn x N/A 49* 24* N/A N/A 33* 40* 62* 78* N/A 53* 42* N/A N/A 67* 73* 16 69* N/A 38* N/A N/A 45* 37* 14 hun Latn x 26 44* 15 40* 44* 62* 76* 34* 39* 23* 31* 30* 43* 50* 20* 61* 38* 39* 50* 38* hye Armn x 34* 18 24 32* 24* 50* 26* 51* 34* 31* 15 40* 49* 31* 56* 51* 33* 31* 45* 41* 43* 29* 16 ibo Latn 34* 43* 18* 45* 46* 49* 52* 52* 64* 30 38* 21* 34 40* 51* 60* 45* 37* 27* 49* 27* 42* 18 58* 55* ilo Latn N/A 46* N/A N/A 53* 55* 61* 75* N/A 40* N/A N/A 57* 60* 52* N/A N/A N/A 15 15 36* 23* ind Latn x 40* 33* N/A 55* 51* 59* 73* 33* 21* 29* N/A 62* 56* 52* N/A isl Latn x N/A 47* 27* 32* 43* 47* 56* 42* 56* N/A 34* 25* 33* 47* 48* 57* 38* N/A 40* 54* 47* 39* 29* 19* ita Latn x N/A 45* 27* 34* 43* 56* 56* 60* 73* N/A 43* 36* 52* 59* 64* 65* 67* N/A 37* 44* 51* 44* 40* jav Latn x 40* N/A N/A 59* 53* 56* 69* 32* 21* N/A N/A 63* 59* 51* N/A N/A jpn Jpan x 27 21 21* 24* 52* 50* 69* 50* N/A N/A N/A N/A N/A N/A N/A N/A N/A 28* 29* 34* 36* 34* kab Latn N/A 40* N/A N/A 35* 37* 36* 47* N/A 56* 24* N/A N/A 58* 56* 32* N/A 15 N/A N/A 47* 49* kac Latn 33* 26 47* 52* 57* 33* 45* 19 14 28* 48* 59* 60* 28* 30* -36* 15 51* 55* kam Latn N/A 46* 20* N/A 55* 56* 58* 59* 70* N/A 32* 26* N/A 55* 56* 66* 20* 51* N/A 18 N/A 31* 34* 24* 52* 43* kan Knda x 24* 19* 22 44* 42* 16 54* 39* 60* 52* 51* 54* 53* 65* 35* 67* 20 24* 14 kas Deva, Arab 29* 64 31* 32* 36* 20* 49* 27* 49* 80* 42* 45* 53* 41* 23 31* 25* 44* 18 kat Geor x 20 28* 51* 48* 50* 49* 39* 29* 43* 45* 44* 44* 21* 56* 34* 31* 53* 32* 17 kaz Cyrl x N/A N/A 25* N/A N/A 45* 39* 51* 69* N/A N/A 30* N/A N/A 72* 72* 27* 50* N/A N/A 29* N/A N/A 15 15 24* kbp Latn N/A 42* 21* N/A N/A 58* 55* 57* 74* N/A 36* 23* N/A N/A 55* 66* 14 48* N/A 27* 30* N/A N/A 44* 22* 41* 50* kea Latn N/A 47* 32* N/A N/A 53* N/A 63* 75* N/A 46* 34* N/A N/A 57* N/A 60* N/A N/A N/A 40* N/A 30* 21* khk Cyrl 18 31* 25 30* 53* 54* 31* 54* 39* 17 28* 35* 64* 65* 45* 21 -33* 19* 18 khm Khmr 38* 40* 33* 38* 49* 52* 47* 67* 39* 40* 26* 33* 60* 64* 43* 23* 26* 17 19* 32* kik Latn N/A 49* 18* 40* 49* 46* 57* 53* 62* N/A 38* 21* 40* 43* 40* 61* 17 45* N/A 25* 27* 24* 54* 41* kin Latn N/A 49* 33* 68* N/A 59* 67* 52* 67* N/A 40* 35* 66* N/A 59* 68* 23* 57* N/A 34* 27* N/A 46* 42* 53* 57* kir Cyrl x 37* 31* 30* N/A N/A 46* 41* 49* 71* 50* 51* 24* N/A N/A 62* 63* 17 57* 41* 24* N/A N/A 15 17 22*
Table 14: Correlations between topic classification results and similarity measures. Continued in next table; full caption in Table 16.
Test language Original Transliterated mBERT lang script   mB. pho inv geo syn gb gen lex chr tri pho inv geo syn gb gen lex chr tri pho inv geo syn gb gen lex chr swt kmb Latn N/A N/A 32* N/A N/A 52* 60* 51* 65* N/A N/A 25* N/A N/A 47* 66* 48* N/A N/A 42* N/A N/A 59* 41* 53* 67* kmr Latn 24* N/A 20* 56* 68* 34* 41* 15 28 N/A 23* 31* 43* N/A 15 16 50* 38* knc Arab, Latn 24 34* 34* 30* 27* 31* 48* 26* 14 38* 39* 52* 52* 35* 27* 17 22 14 17 46* 50* kon Latn N/A N/A 24* 48* N/A 49* N/A 58* 68* N/A N/A 23* 45* N/A 48* N/A 45* N/A N/A 36* N/A 29* N/A 53* 40* kor Hang x 32* 17 20* 23 53* 55* 49* 43* 33* 17 27* 57* 58* 41* 16 23 lao Laoo N/A 30* 26 50 22* 30* 42* 62* N/A 29* 27* 69* 59* 59* 47* N/A 20 -24* 47 18* 22* 22* 23* lij Latn N/A N/A 30* N/A N/A 59* 58* 60* 72* N/A N/A 39* N/A N/A 65* 62* 66* N/A N/A 31* N/A N/A 44* 38* 31* 25* lim Latn N/A N/A 21* N/A N/A 40* 48* 57* 70* N/A N/A 33* N/A N/A 47* 56* 56* N/A N/A 31* N/A N/A 34* 29* 38* 30* lin Latn N/A 45* 20* N/A N/A 46* 51* 55* 65* N/A 39* 20* N/A N/A 46* 53* 17 45* N/A 28* 20* N/A N/A 44* 42* 55* 54* lit Latn x 32* 35* 15 21 29* 42* 63* 77* 46* 34* 54* 43* 48* 57* 32* 65* 46* 54* 54* 47* 37* 14 lmo Latn x N/A N/A 28* N/A 44* 57* 56* 62* 75* N/A N/A 39* N/A 57* 66* 65* 63* N/A N/A 34* N/A 50* 42* 39* 23* 19* ltg Latn N/A N/A 15 N/A N/A 30* N/A 65* 80* N/A N/A 37* N/A N/A 48* N/A 18* 65* N/A N/A 28* N/A N/A 37* N/A 31* 23* ltz Latn x N/A N/A 24* N/A N/A 39* 46* 56* 67* N/A N/A 34* N/A N/A 43* 49* 15 54* N/A N/A 36* N/A N/A 37* 28* 22* lua Latn N/A 25* N/A N/A 46* 48* 55* 68* N/A 21* N/A N/A 41* 45* 50* N/A 26* N/A N/A 40* 39* 57* 55* lug Latn N/A 49* 25* 53* 60* 51* 55* 57* 70* N/A 35* 25* 49* 53* 49* 55* 20* 50* N/A 37* 27* 34* 50* 53* 44* 54* 64* luo Latn 48* 18 34* 60* 39* 32* 56* 66* 35* 17 37* 58* 49* 43* 14 47* 23* 20 40* 55* 55* lus Latn N/A 37* 35* 37* 42* 47* 59* N/A 25* 33* 45* 48* 40* N/A 20 -29* 51* 42* lvs Latn x N/A N/A 20* N/A [—] 31* 45* 65* 80* N/A N/A 44* N/A [47*] 55* 66* 25* 62* N/A N/A 44* N/A [53*] 45* 34* 17 mag Deva N/A N/A 41* N/A 45* 64* N/A 78* 86* N/A N/A 52* N/A 51* 72* N/A 23* 74* N/A N/A N/A 33* N/A 26* 17 mai Deva N/A 35* 41* 64* 56* 65* 61* 74* 87* N/A 55* 53* 50* 57* 73* 74* 16 74* N/A 43 32* 27* 28* 16 mal Mlym x N/A 24* 17 N/A 27* 37* 38* 21* 51* N/A 49* 50* N/A 53* 52* 61* 37* 59* N/A N/A 21 20* 15 mar Deva x N/A 34* 37* 39* 46* 54* 56* 61* 75* N/A 50* 52* 46* 59* 62* 70* 21* 70* N/A 22 29* 32* 24* 23* min Arab, Latn x N/A 35* 46* 56* 56* 41* 60* N/A 33* 22* 42* 64* 60* 53* N/A -18* -23 34* mkd Cyrl x N/A 15 N/A 22 38* 35* 46* 73* N/A 55* 41* N/A 56* 64* 69* 23* 70* N/A 35* N/A 46* 45* 38* 22* 15 mlt Latn N/A 37* 14 N/A 19 59* 72* N/A 26* 28* N/A 35* 16 48* N/A 17 N/A 22 43* 40* mni Beng 22* 24 N/A 46* 49* 48* 28 35* 23* 32* N/A 55* 59* 50* 19 -20* N/A 15 22* 31* mos Latn N/A 43* 18 56* N/A 48* 49* 60* 71* N/A 39* 17 52* N/A 49* 58* 48* N/A 30* 23* N/A 36* 30* 55* 55* mri Latn 40* 39* 26* 40* 40* 47* 46* 44* 58* 35* 32* 29* 34* 42* 53* 53* 36* 29* 21 20 17 16 56* 54* mya Mymr x 24* 15 29* 40* 56* 55* 27* 56* 27 36* 28* 30* 45* 63* 61* 57* nld Latn x N/A 40* 22* 34* 34* 43* 50* 56* 69* N/A 38* 26* 44* 41* 45* 51* 54* N/A 41* 45* 55* 41* 33* nno Latn x N/A N/A 26* N/A N/A 45* 56* 57* 73* N/A N/A 40* N/A N/A 54* 62* 61* N/A N/A 46* N/A N/A 43* 32* nob Latn x 28* 41* N/A 45* 54* 58* 72* 40* 57* N/A 52* 57* 27* 63* 46* 63* N/A 44* 34* npi Deva x N/A N/A 39* N/A 40* 58* 61* 53* 79* N/A N/A 52* N/A 44* 68* 79* 71* N/A N/A N/A 45* 32* 20* 22* nqo Nkoo N/A N/A N/A N/A 52* N/A 22* 57* N/A N/A 15 N/A N/A 50* N/A 15 54* N/A N/A N/A N/A N/A 16 nso Latn N/A N/A 38* N/A N/A 56* 62* 49* 64* N/A N/A 38* N/A N/A 55* 67* 45* N/A N/A 32* N/A N/A 46* 38* 57* 61* nus Latn N/A 35* 33* 47* 52* 52* 25* 41* N/A 21 25 51* 54* 56* 21* N/A 25* 17 21 19* 15 42* 47* nya Latn N/A N/A 36* 40* N/A 57* 61* 57* 69* N/A N/A 40* 41* N/A 61* 66* 21* 51* N/A N/A 28* 28* N/A 49* 47* 55* 58* oci Latn x N/A N/A 27* N/A 43* 59* 60* 60* 74* N/A N/A 42* N/A 57* 67* 67* 66* N/A N/A 35* N/A 48* 41* 35* 16 ory Orya N/A N/A 20* N/A 36* 25* 36* 14 53* N/A N/A 54* N/A 59* 65* 73* 20* 65* N/A N/A N/A 17 26* pag Latn N/A 42* N/A 37* 46* 51* 63* 78* N/A 36* N/A 30* 52* 58* 55* N/A -15 N/A 16 18 39* 28* pan Guru x N/A 23* 17 19 36* 24* 39* 53* N/A 56* 54* 42* 64* 58* 65* 64* N/A 33* 33* 29* 27* 14 pap Latn N/A N/A 39* N/A N/A 46* N/A 62* 73* N/A N/A 39* N/A N/A 52* N/A 23* 58* N/A N/A 23* N/A N/A 40* N/A 35* 27* pbt Arab N/A N/A N/A 21* 37* 40* 58* N/A N/A 31* N/A 29* 29* 51* 15 68* N/A N/A N/A 26* 26* 25* 40* pes Arab x 26 24* 19 33* 33* 47* 49* 75* 34* 43* 42* 50* 42* 60* 23* 83* 34* 18 22 34* 38* 29* 15 plt Latn 36* 24 N/A 54* 46* 32* 41* 31* N/A 53* 48* 33* -17 N/A 16 34* 16 pol Latn x 28 50* 22 30* 29* 59* 74* 24 43* 31* 33* 38* 56* 59* 21* 62* 27 43* 54* 43* 46* 35* 15 por Latn x N/A 47* 31* 35* 46* 62* 59* 60* 74* N/A 42* 40* 49* 59* 66* 63* 66* N/A 31* 40* 49* 39* 34* prs Arab N/A N/A 28* N/A N/A 34* 43* 48* 76* N/A N/A 39* N/A N/A 42* 54* 16 83* N/A N/A 17 N/A N/A 37* 28* 15 quy Latn N/A 36* 42* N/A 19 46* 53* 55* 62* N/A 29* 30* N/A 38* 54* 57* 36* N/A 29* N/A 48* 37* ron Latn x 28 31* 19* 32* N/A 49* 48* 61* 76* 26 31* 37* 50* N/A 58* 54* 25* 65* 24 39* 46* N/A 44* 38* run Latn N/A 43* 32* N/A N/A 57* 67* 50* 64* N/A 32* 33* N/A N/A 56* 68* 20* 58* N/A 37* 35* N/A N/A 62* 56* 54* 68* rus Cyrl x 29* 29* 33* 36* 33* 38* 56* 77* 35* 48* 21* 50* 47* 59* 65* 27* 74* 32* 38* 57* 53* 46* 38* 24* 16 sag Latn 35* 43* 19* 37* N/A 52* 51* 51* 59* 33* 38* 22* 32* N/A 54* 55* 41* 25 28* 17 23 N/A 26* 18 50* 40* san Deva N/A 31* 37* N/A N/A 45* 52* 56* 73* N/A 55* 53* N/A N/A 48* 63* 21* 71* N/A N/A N/A 38* 36* 35* 23* sat Olck N/A 18* 35* 40* 50* 50* 28* 51* N/A 24* 35* 57* 66* 66* 45* 71* N/A 21 21* 16 42* scn Latn x N/A N/A 23* N/A 40* 58* 61* 61* 74* N/A N/A 35* N/A 59* 67* 67* 61* N/A N/A 27* N/A 47* 44* 38* 22* 17 shn Mymr 23 N/A -20* N/A N/A 19* 24* 50* 70* N/A N/A N/A 33* 34* 30* N/A -17 N/A N/A 29* 36* 20* 31* sin Sinh 20 16 N/A N/A 23* 48* 17 52* 55* 45* N/A N/A 40* 58* 19* 50* N/A N/A 17 23* 21* slk Latn x N/A N/A 21* N/A N/A 34* 39* 64* 78* N/A N/A 39* N/A N/A 63* 68* 23* 70* N/A N/A 38* N/A N/A 44* 34* 16 slv Latn x N/A 47* 25* 30 27* 34* 44* 58* 74* N/A 50* 42* 56* 55* 65* 73* 22* 66* N/A 36* 57* 53* 45* 36* 15 smo Latn N/A N/A 14 31* 34* 41* 45* 57* 68* N/A N/A 15 32* 28* 44* 48* 46* N/A N/A 23 19* 24* 59* 56* sna Latn N/A 57* 34* 59* 57* 55* 60* 57* 71* N/A 52* 35* 52* 58* 56* 66* 53* N/A 30* 40* 52* 56* 51* 49* 63* snd Arab 22 25* N/A N/A 28* N/A 43* 62* 33* 45* N/A N/A 34* N/A 28* 74* -24* N/A N/A N/A 31* som Latn 29 50* N/A 50* 66* 29* 41* 43* 53* 16 31* N/A 58* 71* 29* -23 N/A 15 45* 53* sot Latn N/A N/A 44* 60* 60* 63* 51* 63* N/A N/A 35* 56* 55* 66* 17 44* N/A N/A 40* 52* 51* 45* 60* 65* spa Latn x 30 35* 29* 40* N/A 65* 68* 51* 68* 36* 29* 38* 51* N/A 72* 71* 63* 37* 32* 46* N/A 41* 36* srd Latn N/A N/A 28* N/A N/A 54* N/A 58* 73* N/A N/A 37* N/A N/A 64* N/A 63* N/A N/A 31* N/A N/A 46* N/A 30* 27* srp Cyrl x N/A N/A 17 N/A N/A 33* N/A 49* 72* N/A N/A 36* N/A N/A 67* N/A 23* 70* N/A N/A 38* N/A N/A 45* N/A 23* 15 ssw Latn N/A N/A 40* N/A N/A 58* 63* 55* 69* N/A N/A 43* N/A N/A 60* 68* 54* N/A N/A 44* N/A N/A 59* 56* 51* 64* sun Latn x N/A 45* 29* 42* 59* 57* 61* 75* N/A 38* 20* 40* 60* 62* 54* N/A -17 16 swe Latn x N/A 43* 26* 43* 39* 42* 51* 59* 72* N/A 32* 43* 56* 54* 55* 58* 19* 62* N/A 47* 49* 53* 44* 33* swh Latn x 27 51* 31* 46* 55* 54* 56* 57* 67* 24 34* 32* 42* 47* 48* 56* 19* 47* -16 -20 18 szl Latn N/A N/A N/A N/A 32* N/A 56* 70* N/A N/A 24* N/A N/A 57* N/A 22* 68* N/A N/A 37* N/A N/A 44* N/A 25* tam Taml x N/A 22* 19* 26 36* 42* 42* 50* N/A 40* 30* 37* 51* 64* 64* 47* N/A 29* 16 taq Latn, Tfng N/A 30* 20 18* 21* 43* 56* N/A 28* 40* 45* 48* 24* N/A 26* 25* 21 48* 37* tat Cyrl x N/A N/A 30* 39* 28* 48* 42* 59* 77* N/A N/A 36* 45* 47* 64* 64* 24* 61* N/A N/A 41* 34 26* 15 18 24* 14 tel Telu x 25 28* 25 26* 48* 50* 16 50* 40* 68* 47* 35* 48* 53* 65* 64* 28 20 16 tgk Cyrl x N/A N/A 21* 22 N/A 21* 27* 39* 59* N/A N/A 31* 50* N/A 34* 40* 20* 48* N/A N/A 24* N/A 33* 26* 25* tgl Latn x 30 50* 32* 67* 64* 67* 55* 68* 46* 25* 29 67* 69* 73* 54* -21 27 18* tha Thai 22 25* 44* 38* 40* 28* 60* 35* 31* 50* 54* 54* 45* 42* 20* 17 28* tir Ethi N/A 18 N/A N/A 37* 42* 49* 55* N/A 22 24* N/A N/A 29* 33* 31* N/A N/A N/A 15 33* tpi Latn N/A 35* N/A 22 29* N/A 42* 54* N/A 32* N/A 20 24* N/A 39* N/A 26* -14 N/A N/A 53* 50* tsn Latn N/A N/A 38* N/A 61* 57* 65* 46* 59* N/A N/A 37* N/A 59* 55* 66* 42* N/A N/A 34* N/A 41* 45* 42* 56* 64* tso Latn N/A N/A 38* N/A 56* 56* 62* 53* 65* N/A N/A 36* N/A 59* 53* 65* 16 48* N/A N/A 31* N/A 36* 49* 45* 59* 61* tuk Latn N/A N/A N/A 30* 38* 32* 52* 64* N/A N/A 26* N/A 51* 70* 73* 22* 55* N/A N/A N/A 35* 19*
Table 15: Correlations between topic classification results and similarity measures. Continued and explained in next table.
Test language Original Transliterated mBERT lang script mB. pho inv geo syn gb gen lex chr tri pho inv geo syn gb gen lex chr tri pho inv geo syn gb gen lex chr swt tum Latn N/A N/A 42* N/A N/A 67* 74* 48* 64* N/A N/A 42* N/A N/A 63* 76* 53* N/A N/A 25* N/A N/A 47* 45* 56* 58* tur Latn x 25 42* 27* 23* 61* 72* 36* 56* 21* 24 31* 54* 58* 16 52* 35* 33* 20 18 14 19 twi Latn N/A N/A 17 N/A N/A 53* N/A 55* 65* N/A N/A 21* N/A N/A 53* N/A 22* 42* N/A N/A 30* N/A N/A 40* N/A 52* 57* tzm Tfng 26 18 N/A 39* 38* 50* 32* 33* 25* 46* N/A 68* 62* 57* N/A 24* uig Arab N/A N/A 24* 37* N/A 25* 23* 16 58* N/A N/A 22* 44* N/A 74* 74* 17 49* N/A N/A -29* N/A 29* ukr Cyrl x N/A 17 21* 29* 29* 37* 41* 57* 74* N/A 33* 28* 40* 37* 57* 61* 27* 68* N/A 40* 58* 57* 44* 35* 22* 15 umb Latn 40* 46* 26* N/A 48* 50* 57* 48* 60* 32* 46* 30* N/A 47* 50* 61* 15 49* 29 46* 44* N/A 47* 59* 51* 57* 69* urd Arab x N/A 30* 30* 29* N/A 23* 35* 50* 69* N/A 41* 42* 29 N/A 36* 41* 28* 80* N/A 36* N/A 32* 30* uzn Latn x N/A 33* 22 31* 28* 50* 62* N/A 39* 32* 34* 45* 69* 71* 16 61* N/A 33* 25* 20* vec Latn N/A N/A 29* N/A N/A 61* N/A 58* 71* N/A N/A 41* N/A N/A 70* N/A 64* N/A N/A 33* N/A N/A 44* N/A 31* 26* vie Latn x 30 45* 23 33* 53* 59* 39* 50* 43* 23 34* 59* 66* 27* -20 14 -20* war Latn x N/A N/A N/A 42* 63* 64* 58* 71* N/A N/A 24* N/A 48* 67* 69* 54* N/A N/A N/A 15 24* wol Latn 44* 16 21 43* 37* 33* 58* 66* 34* 42* 36* 42* 40* 23* 16 23 20* 52* 42* xho Latn N/A 46* 50* 51* 34* 66* 72* 47* 61* N/A 38* 48* 48* 30* 63* 74* 51* N/A 32* 48* 28 26* 55* 49* 49* 63* ydd Hebr N/A N/A N/A 17 19 42* N/A 24* 17 N/A N/A 30* 30* 16 63* N/A -25* N/A N/A 27* yor Latn x 32* 42* 36* N/A 42* 44* 53* 65* 31* 32* 17 26* N/A 44* 48* 19* 44* N/A 37* 17 yue Hant 43* 42* 56* 55* 69* 30* N/A N/A N/A N/A N/A N/A N/A N/A N/A 44* 38* 28* 32* 39* 35* zsm Latn N/A 43* N/A 40* 53* [44*] 56* 71* N/A 41* 21* N/A 36* 61* [49*] 52* N/A N/A -19 [—] zul Latn 43* 44* 47* 49* N/A 67* 77* 35* 53* 37* 34* 49* 49* N/A 67* 79* 49* 26 35* 42* 34* N/A 54* 52* 52* 65* mean (all) 16 34 17 29 36 45 49 49 65 21 38 29 36 46 55 61 11 55 13 8 17 22 26 25 22 27 21 12;19 31;36 15;19 25;32 33;38 43;46 47;50 46;50 63;66 16;25 35;39 26;30 32;39 43;47 53;57 58;62 9;13 52;56 9;16 6;10 14;20 18;26 22;29 22;27 19;24 24;30 17;24 mean (mbert) 16 33 17 26 32 44 48 49 66 25 40 32 38 46 57 61 14 59 20 0 24 31 32 27 22 13 3 11;21 29;36 14;19 21;30 28;35 41;46 45;50 45;53 63;69 18;30 37;43 29;34 33;42 43;48 54;59 59;63 11;16 56;61 13;25 -1;1 19;27 24;36 24;36 22;30 19;25 10;15 0;4 mean ($\neg$¬mBERT) 16 35 18 32 40 46 50 48 64 17 36 26 34 46 54 60 9 52 6 15 13 14 20 24 22 38 34 10;21 31;37 14;20 26;37 37;43 43;48 47;52 45;50 61;66 10;22 32;38 23;28 28;39 43;48 51;56 57;62 7;11 48;54 1;10 11;18 8;16 8;18 16;25 20;28 18;25 34;41 30;38
Table 16: Correlations (Pearson’s r $\times$ 100, to save space) between topic classification results and similarity measures for each test language, continued. Where a training or test language has multiple datasets (one per writing system), we use language-wise score averages for calculating the correlations. The asterisk* denotes correlations with a p-value below 0.01. Grey cells with a bar (—) denote correlations with a p-value of 0.05 or above. Square brackets [ ] mean that no entry for this ISO code was found in the linguistic databases, so the entry for its macrolanguage code was used instead (aze for azb and azj, apc for ajp, lav for lvs, zlm for zsm). ‘N/A’ means that no correlation score could be calculated due to missing entries in the linguistic databases (or due to missing transliterations in the case of the transliteration experiment). The bottom rows show the mean scores across test languages (‘—’ entries are treated as zeros, ‘N/A’ entries are ignored). Below each row of mean scores are the corresponding 95% confidence intervals (numbers separated by semicolons). ‘avg mBERT’ (‘avg $\neg$¬mBERT’) is for the test languages that are (are not) in mBERT’s pretraining data.